### v2

# Calibrated sensitivity models

Abstract
In causal inference, sensitivity models assess how unmeasured confounders could alter causal analyses, but the sensitivity parameter – which quantifies the degree of unmeasured confounding – is often difficult to interpret. For this reason, researchers sometimes compare the sensitivity parameter to an estimate for measured confounding. This is known as calibration. Although calibration can aid interpretation, it is typically conducted post hoc, and uncertainty in the point estimate for measured confounding is rarely accounted for. To address these limitations, we propose novel calibrated sensitivity models, which directly bound the degree of unmeasured confounding by a multiple of measured confounding. The calibrated sensitivity parameter is interpretable as an intuitive unit-less ratio of unmeasured to measured confounding, and uncertainty due to estimating measured confounding can be incorporated. Incorporating this uncertainty shows causal analyses can be less or more robust to unmeasured confounding than would have been suggested by standard approaches. We develop efficient estimators and inferential methods for bounds on the average treatment effect with three calibrated sensitivity models, establishing parametric efficiency and asymptotic normality under doubly robust style nonparametric conditions. We illustrate our methods with a data analysis of the effect of mothers’ smoking on infant birthweight.
Keywords: causal inference, sensitivity analysis, calibration, doubly robust estimation
1 Introduction
In causal inference, the goal is often to estimate whether exposure to a treatment causes a change in outcomes. Experiments where exposure to treatment is randomized are a classical method to validly estimate causal effects under minimal assumptions, but they are often infeasible, unethical, or too expensive. Therefore, researchers frequently estimate causal effects from observational data, where exposure to treatment is not randomized. To estimate causal effects with observational data, researchers routinely invoke the no unmeasured confounding assumption, which says that the treatment is as-if randomized within observed covariate strata. Unfortunately, this assumption is often implausible, because often there are variables beyond the ones observed that are associated with the treatment and outcomes of the study. Thus, it is imperative to conduct a sensitivity analysis to understand how robust causal analyses are to unmeasured confounding.
We focus on partial identification sensitivity analyses. In simplified terms, these analyses impose a bound
$U\leq\gamma,$ (1)
where $\gamma$ is a sensitivity parameter and $U$ is some quantification of unmeasured confounding, e.g., the difference in counterfactual regression functions (Díaz and van der Laan, 2013) or the odds ratio of the probability of treatment (Rosenbaum, 2002; Tan, 2006). The model in (1) implies bounds on the causal effect which can be estimated from observed data. To understand the impact of unmeasured confounding, one can vary the sensitivity parameter, thereby allowing for different degrees of unmeasured confounding, and estimate bounds on and construct confidence intervals for the causal effect. Often, researchers determine the value of $\gamma$ where the confidence interval includes zero, because it indicates the level of unmeasured confounding where the causal effect estimate is non-significant. If the resulting $\gamma$ value is large, causal analyses are said to be robust to unmeasured confounding; conversely, if the $\gamma$ value is small, analyses are said to be sensitive to unmeasured confounding.
However, it can be difficult to gain intuition for the absolute size of the sensitivity parameter $\gamma$. Therefore, recent research has proposed calibrating (or, benchmarking) results by estimating measured confounding (e.g., Cinelli and Hazlett (2020); Veitch and Zaveri (2020); Franks et al. (2020), among others). Typically, researchers leave out one or more variables from their data, allowing them to act as proxies for unmeasured confounders, and estimate a quantification of measured confounding analogous to the unmeasured confounding $U$ in their sensitivity model (e.g., if $U$ is the odds ratio of propensity scores, then so is measured confounding). Customarily, one can then decide whether the causal effect estimate is robust to unmeasured confounding by comparing the sensitivity parameter to estimated measured confounding. For example, one might compare the level of the sensitivity parameter where confidence intervals for the causal effect include zero to the estimated measured confounding. If the measured confounding were much smaller, this might be evidence that the causal effect estimate is robust to unmeasured confounding, because unmeasured confounders would need to have a larger impact on the causal effect than measured confounders to potentially reverse conclusions from the causal analysis.
This approach, which we refer to as “post hoc calibration”, suffers from two drawbacks. First, uncertainty in the estimate for measured confounding is unaccounted for. As we demonstrate, correctly accounting for this uncertainty can yield different conclusions for how robust the causal effect is to unmeasured confounding — both more and less robust are possible. Second, researchers rarely justify their choice of measured confounding and which covariates are excluded. Therefore, the trade-offs associated with a particular choice of measured confounding are unclear. The choice of measured confounding is important, because it affects how one interprets the robustness of causal analyses.
The lack of interpretability of the sensitivity parameter $\gamma$ and the deficiencies of post hoc calibration motivate our novel approach, the calibrated sensitivity model. In simplified terms, calibrated sensitivity models use sensitivity models as a building block to impose a bound
$U\leq\Gamma M,$ (2)
where $U$ is unmeasured confounding, $M$ is measured confounding, and $\Gamma$ is a sensitivity parameter. In other words, calibrated sensitivity models bound the degree of unmeasured confounding by a multiple of measured confounding.
Calibrated sensitivity models can solve the deficiencies of standard approaches. First, the sensitivity parameter $\Gamma$ is immediately interpretable as a unit-less bound on the ratio of unmeasured confounding divided by measured confounding. Moreover, because measured confounding is included in the model itself, in (2), it is clearer that researchers must explain and justify their choice of measured confounding. Finally, a calibrated sensitivity analysis can be conducted like a standard sensitivity analysis, but in a way that naturally incorporates estimation and uncertainty quantification for measured confounding. And, there is no need for post hoc calibration, because the sensitivity parameter is calibrated by construction.
The goal of this manuscript is to define calibrated sensitivity models, further illustrate their benefits, and demonstrate how to conduct estimation and inference with them. To that end, our primary contributions are threefold. First, we formally define several calibrated sensitivity models. These are based on popular sensitivity models defined on the outcome regression function and propensity score (Luedtke et al., 2015; Rosenbaum, 2002). Second, we establish theory and methods for estimation and inference with calibrated sensitivity models which account for uncertainty in estimating measured confounding. Our estimators attain parametric efficiency and asymptotic normality under doubly-robust style nonparametric conditions on their nuisance function estimators. This allows us to use flexible estimators for nuisance functions like outcome regressions and the propensity score but still conduct valid inference within sensitivity analyses. Finally, we provide an illustrative data analysis, which demonstrates how to analyze data with a calibrated sensitivity model, and highlights the differences between our novel methods and standard approaches. Our code is available at https://github.com/alecmcclean/Calibrated-sensitivity-models
1.1 Structure of the paper
In Section 2, we review standard sensitivity analyses and post hoc calibration approaches for the average treatment effect (ATE), as well as different quantifications of measured confounding and their strengths and weaknesses. In Section 3, we introduce calibrated sensitivity models and provide three examples. In Section 4, we identify bounds on the ATE under the three example models. In Section 5, we develop theory and methods for estimation and inference for bounds on the ATE. In Section 6, we illustrate our methods with a real data analysis on the effect of mothers’ smoking on infant birth weight. In Section 7, we conclude and discuss.
1.2 Notation
[…]
2 Setup and background
2.1 Data, assumptions, and the average treatment effect
We assume we observe $n$ observations $\{Z_{i}\}_{i=1}^{n}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{P}$ where $Z$ is a tuple $(X,A,Y)$, $X\in\mathbb{R}^{d}$ are $d$-dimensional covariates, $A\in\{0,1\}$ is a binary exposure, and $Y\in\mathbb{R}$ is an outcome. We use subscript notation to develop calibrated sensitivity models: $X_{-j}$ excludes the $j^{th}$ covariate, $X_{j}$ is the $j^{th}$ covariate, and, for any set $S\subseteq[d]$, $X_{-S}$ excludes the covariates corresponding to $S$. We let $Z_{-S}=\{X_{-S},A,Y\}$ and assume $\{Z_{-S}\}_{i=1}^{n}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{P}_{-S}$, i.e., the data with only covariates $X_{-S}$ is drawn from some distribution $\mathcal{P}_{-S}$. Additionally, we denote unmeasured confounders by $W$ and define potential outcomes $Y^{a}$ as the outcome that would have been observed under exposure $A=a$. Finally, we will refer to $\pi_{a}(X)=\mathbb{P}(A=a\mid X)$ as the “propensity score” and $\mu_{a}(X)=\mathbb{E}(Y\mid A=a,X)$ as the “outcome regression function”, and generically we will refer to them as “nuisance functions”. Other nuisance functions will be defined when they appear. We make two standard causal assumptions.
Assumption 1.
Consistency: $A=a\implies Y=Y^{a}$.
Assumption 2.
Positivity: $\exists\ \varepsilon>0$ such that $\mathbb{P}\left\{\varepsilon\leq\pi_{1}(X)\leq 1-\varepsilon\right\}=1$.
Consistency says that we observe the potential outcome relevant to the observed exposure. It would be violated if, for example, there were interference between subjects such that one subject’s treatment affected another’s outcome. Positivity says that all subjects have a non-zero probability of exposure. The literature addressing violations of each of these assumptions is too large to summarize here, but see, for example, Tchetgen and VanderWeele (2012) and Westreich and Cole (2010) for discussion of violations of consistency and positivity, respectively.
[…]
$\psi_{-S}=\mathbb{E}\{\mu_{1}(X_{-S})-\mu_{0}(X_{-S})\}.$ (3)
Without no unmeasured confounding or an alternative assumption (e.g., that an instrumental variable or regression discontinuity exists), one cannot identify $\psi_{\ast}$ such that $\psi_{\ast}=\psi$. To address this, there is a large literature on sensitivity analyses, which impose an assumption on the effect of unmeasured confounding.
2.2 Sensitivity analyses
A recent review on sensitivity analysis is Richardson et al. (2014). Modern sensitivity models could arguably be split into two types: models which admit point identification of the causal effect, and models which admit partial (or, set) identification of the causal effect. Generally, point identification models impose a specific form on the relationship between the unmeasured confounder and the observed data. For example, Scharfstein et al. (2021) consider a model of the form $p(Y^{a}=y\mid A=1-a,X=x)=p(Y^{a}=y\mid A=a,X=x)q_{a}(Y^{a}=y,X=x;\gamma)$, where $p$ is the density of $Y^{a}$ and $q_{a}(Y^{a}=y,X=x;\gamma)$ is a researcher-specified transformation which depends on the sensitivity parameter $\gamma$. There are many popular models in the literature (e.g., Robins (1999); Brumback et al. (2004); Scharfstein et al. (2021), among others). Under such an assumption, the ATE (for example) can be identified and estimated, and typically researchers examine how the ATE changes with different values of the sensitivity parameter $\gamma$. We direct interested readers to the literature review mentioned earlier, along with Section 2.2 of Scharfstein et al. (2021).
Partial identification sensitivity analysis, which we focus on in this paper, is the other important type of sensitivity analysis. Generally, a partial identification sensitivity analysis comprises the following steps. First, the researcher imposes a bound $U\leq\gamma$ for $\gamma\in(0,\infty)$ which implies bounds on the causal effect. We denote the lower and upper bounds on the ATE as $\ell:\mathcal{P}\times(0,\infty)\to\mathbb{R}$ and $u:\mathcal{P}\times(0,\infty)\to\mathbb{R}$, respectively. For various levels of $\gamma$, the researcher estimates the bounds and constructs confidence intervals for the bounds and, by extension, the causal effect. Finally, to interpret levels of $\gamma$, the researcher estimates a quantification of measured confounding $M$, and compares levels of $\gamma$ to $\widehat{M}$. Generally, $M$ is a measured quantity analogous to $U$, where one or more variables are left out at a time, capturing measured confounding implied by those variables.
As far as we are aware, these steps constitute a comprehensive sensitivity analysis in the current state of the literature, though the final calibration step, comparing $\gamma$ to $\widehat{M}$, is not always performed. Researchers might also examine other statistics, such as the level of $\gamma$ where confidence intervals for the causal effect include zero, i.e., where a significant effect estimate is nullified.
Within partial identification sensitivity analyses, the key choices are the quantifications of unmeasured confounding $U$ and measured confounding $M$. Here, we review choices of unmeasured confounding, summarizing a non-exhaustive snapshot of the literature, and the next section provides examples of measured confounding. Early partial identification models imposed bounds which follow directly from bounds on the data itself (Manski, 1990; Robins, 1989), e.g., if the outcome is bounded this implies a bound on the ATE. More recent work has bounded the error due to unmeasured confounding at the level of the causal effect itself (Luedtke et al., 2015), the odds ratio of the propensity score (Rosenbaum, 2002; Tan, 2006), the direct change in the propensity score (Masten and Poirier, 2018), the change in or ratio of outcome regression functions (Díaz and van der Laan, 2013; Luedtke et al., 2015), the change in explained variance of the outcome regressions or the propensity score (Cinelli and Hazlett, 2020; Huang and Pimentel, 2022), and the proportion of units confounded (Bonvini and Kennedy, 2022).
Remark 1.
An often over-looked model choice when quantifying unmeasured confounding is the choice of distance, e.g., $L_{p}(\mathbb{P})$ norm. Most models use $L_{\infty}$ norms and bound maximum unmeasured confounding. We consider both maximum unmeasured confounding and average unmeasured confounding ($p=2$) in our examples in Section 3.
2.3 Measured confounding
For both point identification and partial identification sensitivity analyses, researchers have considered calibrating their sensitivity parameters to the observed data by estimating measured confounding (e.g., Franks et al. (2020); Veitch and Zaveri (2020); Cinelli and Hazlett (2020)), and have observed that different quantifications of measured confounding have advantages and disadvantages (e.g., (Cinelli and Hazlett, 2020) Section 4, Veitch and Zaveri (2020) Section 3). In general, there are three choices when constructing measured confounding $M$: the quantification of measured confounding (e.g., effect difference, change in outcome regressions, etc.), what covariate subsets are omitted, and how sub-quantifications of measured confounding across covariate subsets are aggregated.
To quantify measured confounding, the simplest choice is to use a notion of measured confounding that is analogous to the quantification of unmeasured confounding that’s bounded in the sensitivity analysis. For example, if $U=|\psi_{\ast}-\psi|$, then measured confounding would be $|\psi-\psi_{-S}|$ for different sets $S$. However, when the analogous notion of measured confounding is difficult to estimate, researchers may prefer to consider a smooth approximation of it. For example, when unmeasured confounding is an $L_{\infty}$ norm, researchers might prefer to define measured confounding as a smooth approximation of it, because $L_{\infty}$ norms cannot be estimated at $\sqrt{n}$ rates under nonparametric assumptions (Lepski et al., 1999).
The second choice in measured confounding is which subsets are considered. Typically, researchers consider the leave-one-out (LOO) error, which leaves out one covariate at a time; e.g., if one defines measured confounding at the level of the causal effect, the LOO error for the $j^{th}$ covariate is $\left|\psi-\psi_{-j}\right|$, where $\psi_{-j}$ is the functional with variable $X_{j}$ excluded. However, an important disadvantage of the LOO error is that it might be misleading if several variables are similar because it does not capture the joint effects of several similar variables. Therefore, researchers could leave out groups of variables, which we refer to as the leave-some-out (LSO) error. For example, if $S$ is an index referring to some set of covariates (e.g., $S=\{1,2\}$), then the $S$-LSO error is $\left|\psi-\psi_{-S}\right|$. The leave-some-out error can capture the joint effect of several correlated variables that would be missed in the leave-one-out error. Some analyses have also considered leaving out all the covariates or a random half of the covariates (e.g., Rubinstein et al. (2022)).
The third choice is how to aggregate sub-quantifications of measured confounding across the subsets considered. Maximums and averages are natural options. For example, the maximum LOO error at the level of the causal effect is $\max_{j\in[d]}\left|\psi-\psi_{-j}\right|$ while the average leave-one-out error is $\frac{1}{d}\sum_{j=1}^{d}\left|\psi-\psi_{-j}\right|$. Either could be advantageous. Because averages are no larger than maximums, the average error provides tighter bounds than the maximum, which may be appropriate.
3 Calibrated sensitivity analyses
In this section, we outline the generic steps of partial identification calibrated sensitivity analysis, highlighting where they differ from partial identification sensitivity analyses reviewed in Section 2.2, construct three example calibrated sensitivity models, and further explain the advantages of calibrated sensitivity analyses.
Calibrated sensitivity analyses comprise the following steps. First, the researcher imposes a bound $U\leq\Gamma M$ for $\Gamma\in(0,\infty)$ which implies bounds on the causal effect. We denote lower and upper bounds on the ATE as $\mathcal{L}:\mathcal{P}\times(0,\infty)\to\mathbb{R}$ And $\mathcal{U}:\mathcal{P}\times(0,\infty)\to\mathbb{R}$. The bounds differ from a standard sensitivity analysis because $\mathcal{L}$ and $\mathcal{U}$ are different functionals from $\ell$ and $u$ for a standard sensitivity analysis — $\mathcal{L}$ and $\mathcal{U}$ depend on measured confounding. For various levels of $\Gamma$, the researcher estimates the bounds and constructs confidence intervals for the bounds and, by extension, the causal effect. Unlike with standard sensitivity analyses, the method for constructing confidence intervals accounts for uncertainty in estimating measured confounding. These steps constitute a comprehensive calibrated sensitivity analysis. Unlike with standard sensitivity analyses, a final post hoc calibration step is not necessary, because $\Gamma$ is already calibrated by construction.
3.1 Three example calibrated sensitivity models
The first model we consider is the maximum leave-one-out effect differences model (which we also refer to as the “effect differences model”).
Calibrated sensitivity model 1.
(Maximum leave-one-out effect differences)
$\left|\psi_{\ast}-\psi\right|\leq\Gamma\max_{j\in[d]}\left|\psi-\psi_{-j}\right|,$ (4)
where $\Gamma\in(0,\infty)$ is the sensitivity parameter, $\psi$ is the adjusted mean difference in (3) with $X$ covariates (i.e., all covariates) and $\psi_{-j}$ is the adjusted mean difference with $X_{-j}$ covariates (i.e., without covariate $X_{j}$).
The maximum LOO effect differences model uses the maximum leave-one-out error and quantifies unmeasured and measured confounding at the level of the causal effect. This quantification of unmeasured confounding was considered in Luedtke et al. (2015). While quantifying confounding at the level of the causal effect is less popular than the other approaches we consider in the next two examples, we include this approach because the resulting analysis is intuitive and simple, and therefore eases exposition of the main ideas in this paper. Furthermore, this choice of $U$ and $M$ has other benefits when considered within a calibrated sensitivity model. First, this model generalizes to any causal effect which is identifiable under no unmeasured confounding. Moreover, estimation is straightforward, and does not require the derivation of new estimators; instead, one can re-run an estimator for $\psi$, but with fewer covariates included.
The second model is the maximum leave-one-out $L_{\infty}$ propensity score odds ratio model (which we also refer to as the “odds ratio model”).
Calibrated sensitivity model 2.
(Maximum leave-one-out $L_{\infty}$ propensity score odds ratio)
$\sup_{x,w,\widetilde{w}}\left|\log\left[\frac{\text{odds}\{\pi_{1}(x,w)\}}{\text{odds}\{\pi_{1}(x,\widetilde{w})\}}\right]\right|\leq\Gamma\max_{j\in[d]}\sup_{x_{-j},x_{j},\widetilde{x}_{j}}\left|\log\left[\frac{\text{odds}\{\pi_{1}(x_{-j},x_{j})\}}{\text{odds}\{\pi_{1}(x_{-j},\widetilde{x}_{j})\}}\right]\right|$ (5)
where $\Gamma\in(0,\infty)$ is the sensitivity parameter, $\pi_{1}$ is the propensity score, and $\text{odds}(p)=p/(1-p)$.
The maximum LOO odds ratio model uses the maximum leave-one-out error again and a popular quantification of unmeasured confounding — the maximum propensity score odds ratio induced by the unmeasured confounder (Rosenbaum, 2002; Yadlowsky et al., 2022). We state the model in terms of the log-odds ratio so that the calibrated sensitivity parameter $\Gamma$ takes the same range and has the same meaning as in the other two models. In the odds ratio model, the unmeasured and measured confounding are defined in terms of a supremum, which is a popular definition in the literature. A drawback of this is that it is necessary to estimate the supremum, and $\sqrt{n}$-estimation is not possible under nonparametric assumptions (Lepski et al., 1999). Therefore, in Section 5, we will consider the projection of measured confounding onto a finite-dimensional model, for which one can construct $\sqrt{n}$-efficient estimators and methods for inference.
The third model is the average leave-some-out $L_{2}$ outcome regression differences model (which we also refer to as the “outcome model”).
Calibrated sensitivity model 3.
(Average leave-some-out $L_{2}$ outcome regression differences) For $a\in\{0,1\}$,
$\left\lVert\mathbb{E}(Y^{a}\mid A=a,X)-\mathbb{E}(Y^{a}\mid A=1-a,X)\right\rVert_{2}^{2}$
$\hskip 72.26999pt\leq\frac{\Gamma}{|\mathcal{S}|}\sum_{S\in\mathcal{S}}\left\lVert\mu_{a}(X_{-S})-\mathbb{E}\big{\{}\mu_{a}(X)\mid A=1-a,X_{-S}\big{\}}\right\rVert_{2}^{2},$ (6)
where $\Gamma\in(0,\infty)$ is the sensitivity parameter and $\mathcal{S}$ is a set of subsets of $[d]$.
The average LSO outcome model is also based on a popular quantification of unmeasured confounding — the difference between the potential outcome regressions under the observed treatment and the opposite treatment (Díaz and van der Laan, 2013; Luedtke et al., 2015). We use this model to demonstrate alternative choices one might make when constructing a calibrated sensitivity model. First, unmeasured and measured confounding are in terms of $L_{2}^{2}$ norms rather than the usual $L_{\infty}$ norm. Second, measured confounding is the average LSO error rather than the maximum LOO error in the two previous models.
We conclude this section with a general assumption. It asserts that the overall quantification of measured confounding is bounded and non-zero.
Assumption 3.
Bounded and non-zero measured confounding:
  1. 1.
    Under the effect differences model, $|\psi-\psi_{-j}|<\infty$ for all $j\in[d]$ and there exists $j\in[d]$ such that $|\psi-\psi_{-j}|>0$.
  2. 2.
    Under the odds ratio model, $\sup_{x_{-j},x_{j},\widetilde{x}_{j}}\left|\log\left[\frac{\text{odds}\{\pi_{1}(x_{-j},x_{j})\}}{\text{odds}\{\pi_{1}(x_{-j},\widetilde{x}_{j})\}}\right]\right|<\infty$ for all $j\in[d]$ and there exists $j\in[d]$ such that $\sup_{x_{-j},x_{j},\widetilde{x}_{j}}\left|\log\left[\frac{\text{odds}\{\pi_{1}(x_{-j},x_{j})\}}{\text{odds}\{\pi_{1}(x_{-j},\widetilde{x}_{j})\}}\right]\right|>0$.
  3. 3.
    Under the outcome model, for $a\in\{0,1\}$, $\lVert\mu_{a}(X_{-S})-\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-s}\}\rVert_{2}^{2}<\infty$ for all $S\in\mathcal{S}$ and there exists $S\in\mathcal{S}$ such that $\lVert\mu_{a}(X_{-S})-\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-s}\}\rVert_{2}^{2}>0$.
This assumption is very mild, but is necessary so that the calibrated sensitivity analyses are meaningful. If measured confounding is infinite, then the bound on unmeasured confounding is also infinite, and the bounds on the ATE will be infinitely wide. Meanwhile, if measured confounding is zero, then the bound on unmeasured confounding is again zero, and the model assumes that no unmeasured confounding holds.
3.2 Advantages of calibrated sensitivity models
The three examples illustrate how calibrated sensitivity models have several advantages, compared to standard sensitivity models. First, unlike the standard sensitivity parameter $\gamma$, which researchers often struggle to interpret, the calibrated sensitivity parameter $\Gamma$ is calibrated by construction and is an intuitive, unit-less bound on the ratio of unmeasured to measured confounding. Second, with calibrated sensitivity models, confidence intervals for the causal effect that account for the uncertainty in estimated measured confounding can be constructed. While this is possible with sensitivity models and post hoc calibration, it has not previously been considered in the literature, and it is simpler with calibrated sensitivity models. We provide methods to do so subsequently, and demonstrate how results can change with a data analysis in Section 6. Third, incorporating measured confounding as an explicit assumption in a calibrated sensitivity model can encourage researchers to better justify their choice of measured confounding.
Moreover, calibrated sensitivity models offer a further potential benefit, beyond solving the issues inherent in standard approaches. Namely, calibrated sensitivity models can be constructed with less popular but perhaps more useful quantifications of unmeasured confounding, which might increase the adoption of sensitivity analyses in practice. For instance, the effect differences model (in calibrated sensitivity model 1), which bounds unmeasured confounding at the level of the causal effect, is not widely used for standard sensitivity analyses, presumably due interpretability issues for the bound $|\psi_{\ast}-\psi|\leq\gamma$. However, within a calibrated sensitivity model, it is more interpretable, and it could offer advantages over standard quantifications of unmeasured confounding. Indeed, the effect differences model is very flexible, as it can be applied to sensitivity analyses with any causal effect identified under unmeasured confounding. By contrast, many popular quantifications of unmeasured confounding are not so flexible; for example, models defined on the odds ratio of propensity scores do not easily generalize beyond binary treatment (Rosenbaum, 2002; Tan, 2006). Other quantifications of unmeasured confounding, such as distributional distances or average strengths of unmeasured confounding (Luedtke et al., 2015; Zhang and Zhao, 2022; Jin et al., 2022), can also be interpretable within a calibrated sensitivity framework, and may offer advantages over quantifications of unmeasured confounding which are currently more popular.
4 Partial identification
In this section, we provide partial identification results for bounds on the ATE from the three examples in Section 3, and then establish that the bounds are differentiable with respect to measured confounding, which we use to provide estimation convergence guarantees subsequently.
Proposition 1.
(Partial identification) Suppose Assumptions 1 and 2 hold. Under the maximum LOO effect differences model (calibrated sensitivity model 1),
$\mathcal{U}(\Gamma)$ $=\psi+\Gamma\max_{j\in[d]}\left|\psi-\psi_{-j}\right|\text{ and }$
$\mathcal{L}(\Gamma)$ $=\psi-\Gamma\max_{j\in[d]}\left|\psi-\psi_{-j}\right|$ (7)
Under the maximum LOO odds ratio model (calibrated sensitivity model 2),
$\mathcal{L}(\Gamma)$ $=\mathbb{E}\Big{[}(1-A)\theta_{1}^{-}\{X;\exp(\Gamma M)\}-A\theta_{0}^{+}\{X;\exp(\Gamma M)\}\Big{]}\text{ and }$
$\mathcal{U}(\Gamma)$ $=\mathbb{E}\Big{[}(1-A)\theta_{1}^{+}\{X;\exp(\Gamma M)\}-A\theta_{0}^{-}\{X;\exp(\Gamma M)\}\Big{]}$ (8)
where, e.g.,
$\theta_{1}^{-}(X;t)=\inf\Big{\{}\mathbb{E}\{YL(Y)\mid A=1,X\}:L\in\mathbb{L}\Big{\}},$ (9)
and $\mathbb{L}=\left\{L:\mathbb{R}\to\mathbb{R}\text{ measurable }:\begin{array}[]{l}0\leq L(y)\leq tL(\tilde{y})\text{ for all }y,\tilde{y},\\ \mathbb{E}\{L(Y)\mid A=1,X\}=1\end{array}\right\}$, and
$M=\max_{j\in[d]}\sup_{x_{-j},x_{j}}\left|\log\left[\frac{\text{odds}\{\pi_{1}(x_{-j},x_{j})\}}{\text{odds}\{\pi_{1}(x_{-j},\widetilde{x}_{j})\}}\right]\right|.$ (10)
Finally, under the average LSO outcome model (calibrated sensitivity model 3),
$\mathcal{U}(\Gamma)$ $=\psi+\Gamma\sum_{a\in\{0,1\}}\lVert\pi_{1-a}(X)\rVert_{2}\sqrt{\frac{1}{|\mathcal{S}|}\sum_{S\in\mathcal{S}}\lVert\mu_{a}(X_{-S})-\mathbb{E}\big{\{}\mu_{a}(X)\mid A=1-a,X_{-S}\}\rVert_{2}^{2}}\text{ and }$
$\mathcal{L}(\Gamma)$ $=\psi-\Gamma\sum_{a\in\{0,1\}}\lVert\pi_{1-a}(X)\rVert_{2}\sqrt{\frac{1}{|\mathcal{S}|}\sum_{S\in\mathcal{S}}\lVert\mu_{a}(X_{-S})-\mathbb{E}\big{\{}\mu_{a}(X)\mid A=1-a,X_{-S}\}\rVert_{2}^{2}}$ (11)
Proposition 1 shows that all three calibrated sensitivity models induce bounds on the ATE which follow naturally from the bounds induced by the relevant sensitivity model, but including measured confounding in the bounds. Expression (7) follows directly from the definition of the model; (8) follows from the definition of the model and Lemma 2.1 in Yadlowsky et al. (2022); and (11) follows by Hölder’s inequality. A formal proof, and all subsequent proofs, are delayed to the supplementary materials. For partial identification in the odds ratio model, only $\theta_{1}^{-}$ is defined, in (9). Taking a supremum instead of an infimum gives $\theta_{1}^{+}$, and swapping the conditioning to $A=0$ from $A=1$ gives $\theta_{0}^{-}$ and $\theta_{0}^{+}$.
4.1 Differentiable bounds
When estimating the bounds, we will rely on them being differentiable with respect to measured confounding to allow use of Taylor’s theorem and the delta method when providing convergence guarantees. Under Assumption 3, this is trivially true for the effect differences and outcome regression models because measured confounding appears linearly in the bounds, but proving this is more involved for the odds ratio model. The derivative for the upper bound is established in the next result, while the derivative for the lower bound follows by a similar analysis. In addition to Assumption 3, the result relies on the positivity assumption (Assumption 2) and a regularity assumption (Assumption 7), detailed in the supplementary materials, which asserts that the covariate density, propensity score, and $\theta_{a}^{\pm}$ from (9) are continuous in $x$, while the outcome $Y$ is bounded and has continuous and upper bounded conditional density.
Lemma 1.
(Differentiable and monotone bounds in $M$ for the odds ratio model)
Let the upper bound $\mathcal{U}(\Gamma)$ be defined as in (8) and $M$ denote measured confounding as in (10). Suppose Assumptions 2, 3, and 7 hold. Then,
$\frac{\partial}{\partial M}\mathcal{U}(\Gamma)=\Gamma\mathbb{E}\left\{\pi_{0}(X)\frac{\widetilde{f}_{1}(X;\theta_{1}^{+})}{\nu_{1}^{+}(X)}\right\}+\Gamma\exp(\Gamma M)\mathbb{E}\left\{\pi_{1}(X)\frac{f_{0}(X;\theta_{0}^{-})}{\nu_{0}^{-}(X)}\right\},$ (12)
where $f_{a},\widetilde{f}_{a}$, and $\nu_{a}^{\pm}$ are defined in Lemma 2 in the supplementary materials. Moreover,
$\exists\ C>0\text{ such that }-C<\frac{\partial}{\partial M}\mathcal{L}(\Gamma)<0<\frac{\partial}{\partial M}\mathcal{U}(\Gamma)<C.$
Lemma 1 shows that the bounds implied by the odds ratio model widen as measured confounding increases. Unlike the effect differences and outcome models, the bounds in the odds ratio model widen in a way that’s non-linear with $M$.
5 Estimation and inference
In this section, we establish convergence guarantees for estimators of measured confounding and bounds on the ATE with two of the three example models. In the supplementary materials, we establish estimation with the outcome model and demonstrate when sensitivity analyses with post hoc calibration might over- or under-estimate robustness to unmeasured confounding by not accounting for uncertainty in estimating measured confounding.
Before stating the results, we make several general observations. First, many of the convergence guarantees in this paper are doubly robust, showing $\sqrt{n}$-estimation can be possible even when nuisance functions are estimated at slower rates. This occurs because the estimators are based on the efficient influence function (EIF), which allows their bias to be a second-order product of errors, so they can achieve $\sqrt{n}$-efficiency even when the nuisance functions are estimated at slower-than-$\sqrt{n}$ rates (Chernozhukov et al., 2018; Kennedy, 2022). These slower rates could be achieved under nonparametric structural assumptions on the nuisance functions, such as smoothness, sparsity, or bounded variation (Györfi et al., 2002).
Second, the results require stronger convergence guarantees on the nuisance function estimators than what is required for $\sqrt{n}$-consistent estimation of the bounds on the ATE in a standard sensitivity analysis. This occurs because it is necessary to construct $\sqrt{n}$-consistent estimators for measured confounding, which require accurate nuisance function estimators with multiple sets of covariates. By contrast, a standard sensitivity analysis only requires accurate nuisance function estimators with all covariates. However, it is worth noting that the requirements here are no stronger than what would be required to construct $\sqrt{n}$-consistent estimators of measured confounding in a typical post hoc calibration analysis.
Third, throughout, we assume the nuisance functions estimators are constructed on a sample of $n$ observations which is separate and independent from the sample used to estimate measured confounding and the bounds on the ATE. Sample splitting allows us to avoid imposing Donsker or stability conditions on the nuisance function estimators (e.g., van der Vaart and Wellner (1996); Chen et al. (2022)). Although this seemingly cuts the sample size in half, one could retain full sample efficiency by swapping folds, repeating the estimator, and averaging the two estimates. For better stability, one could use more than two splits; five and ten are common. To simplify notation, we focus on the single split estimator.
Finally, the causal consistency assumption (Assumption 1) is not required for these results until Section 5.3, when we construct confidence intervals for the ATE. This is because Theorems 1 & 2 focus on estimating the upper bound $\mathcal{U}(\Gamma)$, which is a non-causal functional of the observed data. By contrast, the positivity assumption (Assumption 2) is required as a regularity assumption so that $\mu_{a}(X)$ is well-defined for $a\in\{0,1\}$.
5.1 Effect differences
First, we consider estimating the bound on the ATE in the effect differences model (calibrated sensitivity model 1).
Definition 1.
Let $\mathcal{U}(\Gamma)$ be as in (7), and construct an estimator for the upper bound as
$\widehat{\mathcal{U}}(\Gamma):=\mathbb{P}_{n}\{\widehat{\phi}(Z)\}+\Gamma\max_{j\in[d]}\left|\mathbb{P}_{n}\{\widehat{\phi}(Z)-\widehat{\phi}(Z_{-j})\}\right|,$ (13)
where $\phi(Z_{-j})$ is the EIF of the adjusted mean difference $\psi_{-j}$, i.e.,
$\phi(Z_{-j})=\mu_{1}(X_{-j})-\mu_{0}(X_{-j})+\left\{\frac{A}{\pi_{1}(X_{-j})}-\frac{1-A}{\pi_{0}(X_{-j})}\right\}\{Y-\mu_{A}(X_{-j})\},$ (14)
and the estimated nuisance functions constituting $\widehat{\phi}$ are constructed on a separate sample ($\{\widehat{\pi}_{a},\widehat{\mu}_{a}\}$ for $a\in\{0,1\}$ and $X_{-j}$ for $j\in[d]\cup\emptyset$).
The upper bound is identified as $\mathcal{U}(\Gamma)=\psi+\Gamma\max_{j\in[d]}|\psi-\psi_{-j}|$. Therefore, the estimator $\widehat{\mathcal{U}}(\Gamma)$ plugs in estimators for each adjusted mean difference and takes the maximum difference $\left|\widehat{\psi}-\widehat{\psi}_{-j}\right|$ across $j\in[d]$. The estimator uses the EIF of the adjusted mean difference, which is well-studied in the doubly robust estimation literature (Robins et al., 1994). As a result, the convergence guarantees will be doubly robust, as communicated in Theorem 1 below. To facilitate straightforward $\sqrt{n}$-inference when measured confounding is a maximum, we invoke a separation condition.
Assumption 4.
Separation of maximum in the effect differences model: There exists $j^{\prime}\in[d]$ such that $\left|\psi-\psi_{-j^{\prime}}\right|>\left|\psi-\psi_{-j}\right|\ \forall\ j\in[d]\setminus j^{\prime}$.
This assumption is mild, but could be relaxed by considering a smooth approximation of the maximum, such as the LogSumExp function. With the separation condition in hand, the next result establishes convergence guarantees for the estimator of the upper bound on the ATE under the effect differences model.
Theorem 1.
(Maximum LOO effect differences model) Let $\widehat{\mathcal{U}}(\Gamma)$ and $\mathcal{U}(\Gamma)$ be as in Definition 1, and let $j^{\prime}=\operatorname*{arg\,max}_{j\in[d]}\left|\psi-\psi_{-j}\right|.$ Suppose Assumptions 2-4 hold and
  1. 1.
    for $j\in[d]\cup\emptyset$, $\widehat{\pi}_{1}(X_{-j})$ is bounded away from zero and one
  2. 2.
    for $j\in[d]\cup\emptyset$, $\widehat{\phi}$ is consistent in the sense that $\lVert\widehat{\phi}(Z_{-j})-\phi(Z_{-j})\rVert_{2}\stackrel{{\scriptstyle p}}{{\rightarrow}}0$, and
  3. 3.
    for $j\in\{j^{\prime},\emptyset\}$, $\lVert\widehat{\pi}_{1}(X_{-j})-\pi_{1}(X_{-j})\rVert_{2}\left(\sum_{a\in\{0,1\}}\lVert\widehat{\mu}_{a}(X_{-j})-\mu_{a}(X_{-j})\rVert_{2}\right)=o_{\mathbb{P}}(n^{-1/2})$.
Then,
$\widehat{\mathcal{U}}(\Gamma)-\mathcal{U}(\Gamma)=(\mathbb{P}_{n}-\mathbb{E})\Big{[}\phi(Z)+\text{sign}(\psi-\psi_{-j^{\prime}})\Gamma\big{\{}\phi(Z)-\phi(Z_{-j^{\prime}})\big{\}}\Big{]}+o_{\mathbb{P}}(n^{-1/2}).$ (15)
Theorem 1 establishes that the error of the estimator for the upper bound behaves like a centered sample average plus asymptotically negligible error under doubly robust conditions on the nuisance function estimators. Indeed, because the estimator uses the EIF of the adjusted mean difference, its error follows the form of the doubly robust estimator for the adjusted mean difference, requiring that the propensity score and outcome regression are estimated at a $\sqrt{n}$-rate in product.
Here, we give some intuition for the conditions of the result. Condition 1 is necessary so that the bias of each adjusted mean difference estimator is bounded, while condition 2 is a weak consistency condition necessary for controlling the empirical process terms in the error of each adjusted mean difference estimator. Crucially, condition 3 is necessary to control the bias of $\widehat{\mathcal{U}}(\Gamma)$. Interestingly, condition 3 only requires that the bias of the estimator for maximum measured confounding ($|\psi-\psi_{-j^{\prime}}|$) converges at a $\sqrt{n}$-rate, while the non-maximums can be estimated consistently at any rate (which is guaranteed by condition 2). Intuitively, this occurs because the estimators for the non-maximums are only used to find the index of the true maximum, which is much easier statistically than estimating the value of the maximum. Indeed, access to merely consistent estimators of $|\psi-\psi_{-j}|$ for all $j\in[d]$, including the maximum, is enough to guarantee that the estimator for the maximum index, $\widehat{j}$, converges arbitrarily quickly to the true maximum index, $j^{\prime}$. A technical lemma establishing this is provided in Appendix G.
One can gain intuition for how the confidence interval for the ATE can be wider or narrower with a calibrated sensitivity model compared to a standard sensitivity model by examining (15). The limiting variance will determine the size of an asymptotically valid confidence interval. Without accounting for uncertainty in estimating measured confounding, as in a standard sensitivity analysis, the limiting variance is $\mathbb{V}\{\phi(Z)\}$. By contrast, (15) yields the limiting variance $\mathbb{V}\left[\phi(Z)+\text{sign}(\psi-\psi_{-j^{\prime}})\Gamma\big{\{}\phi(Z)-\phi(Z_{-j^{\prime}})\big{\}}\right]$. This could be larger or smaller than $\mathbb{V}\{\phi(Z)\}$, depending on the covariance between $\phi(Z)$ and $\text{sign}(\psi-\psi_{-j^{\prime}})\Gamma\big{\{}\phi(Z)-\phi(Z_{-j^{\prime}})\big{\}}$. We provide further analysis in the supplementary materials. Similar principles apply with other models, but with more complicated analysis.
5.2 Odds ratio
Next, we consider estimating the upper bound on the ATE in the odds ratio model (calibrated sensitivity model 2). There are two extra nuances to this estimator. First, because the quantification of measured confounding is a supremum, it is not possible to estimate it at a $\sqrt{n}$-rate under nonparametric assumptions (Lepski et al., 1999). Instead, we target the best projection of the propensity score onto a finite-dimensional model, and use that to estimate measured confounding. This approach has a long history in statistics (e.g., Huber et al. (1967)) and has been studied in many contexts in causal inference (e.g., Semenova and Chernozhukov (2021); Kennedy et al. (2023)). Second, the nuisance functions depend on measured confounding (see, Proposition 1). Therefore, the estimator for the upper bound estimates measured confounding within the training sample, and uses it as an input to estimate the nuisance functions for the upper bound. Finally, it estimates the bound in a separate estimation sample, as with the effect differences model.
For ease of exposition, we impose a further mild assumption on the covariates.
Assumption 5.
Bounded covariates: The support of the covariates is the d-dimensional unit cube.
This allows for construction of an estimator for measured confounding by finding the supremum over the unit cube. This assumption could be relaxed to any known and bounded support. Future work could relax this assumption entirely and incorporate an estimator of the support of the covariates.
Definitions 2 and 3 below define the estimator for measured confounding and the estimator for the upper bound, respectively. We split the estimator into two parts to clearly illustrate how the estimator for the best projection of measured confounding is constructed. Estimating the best projection of the propensity score corresponds to maximum likelihood estimation with binary regression (see, e.g., van der Vaart (2000), Example 5.40). Definition 2 provides a specific example of binary regression — logistic regression with no interactions — but the ideas generalize to other link functions and to transformations of the covariates.
Definition 2.
Suppose Assumption 5 holds. Let $\Psi(x)=\frac{1}{1+\exp(-x)}$ denote the logistic function. Define $\pi_{1}^{\perp}(X)=\Psi(\beta^{T}X)$, where $\beta=\operatorname*{arg\,max}_{b\in B}\mathbb{E}\left[\Psi(b^{T}X)^{A}\left\{1-\Psi(b^{T}X)\right\}^{1-A}\right]$, where $B\subset\mathbb{R}^{d}$ is the parameter space. Construct an estimator for the propensity score as $\widehat{\pi}_{1}^{\perp}=\Psi(\widehat{\beta}^{T}X)$ where $\widehat{\beta}=\operatorname*{arg\,max}_{b\in B}\mathbb{P}_{n}\left[\Psi(b^{T}X)^{A}\left\{1-\Psi(b^{T}X)\right\}^{1-A}\right]$. Because the model uses the logistic link, contains no interactions, and the support of the covariates is the unit cube, measured confounding $M$ (defined in (10)) corresponds to the largest absolute coefficient of $\beta$, i.e., $M=\max_{j\in[d]}|\beta_{j}|$, where $\beta_{j}$ is the $j^{th}$ coefficient of $\beta$. Therefore, let $\widehat{M}=\max_{j\in[d]}|\widehat{\beta}_{j}|.$
As with the effect differences model, to facilitate $\sqrt{n}$-inference when measured confounding is a maximum, we invoke a separation condition.
Assumption 6.
Separation of maximum in the odds ratio model: There exists $j^{\prime}\in[d]$ such that $\left|\beta_{j^{\prime}}\right|>\left|\beta_{j}\right|\ \forall\ j\in[d]\setminus j^{\prime}$.
Definition 3.
Let $M$ and $\widehat{M}$ be as in Definition 2 and $\mathcal{U}(\Gamma)$ as in (8). Construct an estimator for the upper bound as $\widehat{\mathcal{U}}(\Gamma):=\mathbb{P}_{n}\big{[}\varphi_{U}\{Z;\widehat{\eta}(\Gamma\widehat{M})\}\big{]}$, where
  • $\varphi_{U}$ is the EIF for the upper bound on the ATE given in Yadlowsky et al. (2022), also defined in (44) in the supplementary materials,
  • the estimated nuisance functions $\widehat{\eta}=\{\widehat{\theta}_{a}^{\pm},\widehat{\nu}_{a}^{\pm},\widehat{\pi}_{a}\}$ for $a\in\{0,1\}$ are constructed on a separate sample, where $\theta_{a}^{\pm}$ are defined in Proposition 1 and $\nu_{a}^{\pm}$ are defined in Lemma 2 in the supplementary materials, and
  • $\widehat{M}$ is constructed on a separate sample, according to Definition 2.
The next result provides a convergence guarantee for the estimator in Definition 3.
Theorem 2.
(Maximum LOO odds ratio model) Let $\mathcal{U}(\Gamma)$, $\widehat{\mathcal{U}}(\Gamma)$, $\varphi_{\mathcal{U}}$, and $\widehat{\eta}$ be as in Definition 3, and let $j^{\prime}=\operatorname*{arg\,max}_{j\in[d]}|\beta_{j}|$. Suppose Assumptions 2-3 and 5-7 hold, and
  1. 1.
    the distribution of $X$ is not concentrated on a $(d-1)$-dimensional affine subspace of its support,
  2. 2.
    $\beta$ is at an inner point of $B$, the set of possible parameter values,
  3. 3.
    $\widehat{\eta}(\Gamma\widehat{M})$ is consistent for $\eta(\Gamma M)$ in the sense that $\left\lVert\varphi_{U}\{Z;\widehat{\eta}(\Gamma\widehat{M})\}-\varphi_{U}\{Z;\eta(\Gamma M)\}\right\rVert_{2}\stackrel{{\scriptstyle p}}{{\rightarrow}}0$, and
  4. 4.
    the nuisance function estimators satisfy $\mathbb{P}\Big{[}\varphi_{U}\{Z;\widehat{\eta}(\Gamma\widehat{M})\}-\varphi_{U}\{Z;\eta(\Gamma\widehat{M})\}\Big{]}=o_{\mathbb{P}}(n^{-1/2})$, where $\mathbb{P}$ denotes expectation conditional on the training data.
Then,
$\widehat{\mathcal{U}}(\Gamma)-\mathcal{U}(\Gamma)=(\mathbb{P}_{n}-\mathbb{E})\left[\varphi_{U}\{Z;\eta(\Gamma M)\}+\left\{\frac{\partial}{\partial M}\mathcal{U}(\Gamma)\right\}\phi_{M}(Z)\right]+o_{\mathbb{P}}(n^{-1/2}),$ (16)
where the derivative $\frac{\partial}{\partial M}\mathcal{U}(\Gamma)$ is defined in (12), and $\phi_{M}(Z)=e_{j^{\prime}}^{T}I_{\beta}^{-1}\ \big{\{}\text{sign}(\beta_{j^{\prime}})s(Z;\beta)\big{\}}$, where $e_{j}$ is the $j^{th}$ unit vector, $I$ is the Fisher information, and $s(\cdot)$ is the score function. With the model $\Psi(\beta^{T}X)$ considered here, $I_{\beta}=\mathbb{E}\left[\frac{\Psi^{\prime}(\beta^{T}X)^{2}}{\Psi(\beta^{T}X)\{\beta^{T}X-\Psi(\beta^{T}X)\}}XX^{T}\right]$ and $s(Z;\beta)=\frac{A-\Psi(\beta^{T}X)}{\Psi(\beta^{T}X)\{\beta^{T}X-\Psi(\beta^{T}X)\}}\Psi(\beta^{T}X)X.$
Theorem 2 shows that the estimator for the upper bound in the odds ratio model behaves like a centered sample average plus asymptotically negligible error. Here, we give some intuition for the conditions of the result. Conditions 1 and 2 are necessary so that the estimator $\widehat{M}$ successfully targets $M$, the best projection of measured confounding onto the logistic model with no interactions. Condition 1 is necessary so that the parameter $\beta$ is identifiable, in the statistical non-causal sense; see, e.g., pg. 62 in van der Vaart (2000). Indeed, if this condition does not hold, then infinite parameter estimates can exist (see, e.g., Section 5.4.2., Agresti (2015)). Condition 2 is a standard condition necessary for establishing the convergence guarantees of M-estimators (see van der Vaart (2000), Section 5), of which maximum likelihood estimation, used to construct $\widehat{\beta}$, is an example. Conditions 3 and 4 control the error of the estimator of the upper bound. Condition 3 is a mild consistency condition on the nuisance function estimators, which allows us to control the empirical process term. Meanwhile, condition 4 asserts that the conditional bias due to estimating the nuisance functions — with estimated measured confounding $\widehat{M}$ as an input — is asymptotically negligible. This allows us to establish that the bias of the estimator $\widehat{\mathcal{U}}(\Gamma)$ is asymptotically negligible. This is the same doubly robust-style condition from Yadlowsky et al. (2022) Assumption A.2, but with estimated measured confounding $\widehat{M}$ multiplied by the calibrated sensitivity parameter $\Gamma$ rather than a fixed sensitivity parameter $\gamma$. We anticipate this holds with many nonparametric estimators under assumptions like smoothness or sparsity.
5.3 Inference
The results in the previous sections establish that estimators for the bounds on the ATE which incorporate an estimator for measured confounding can be $\sqrt{n}$-efficient and asymptotically linear under nonparametric doubly-robust style conditions. By the central limit theorem, Wald-type confidence intervals can be constructed using the normal distribution and the influence functions in Theorems 1 & 2 (here, we mean influence functions in the sense of asymptotically linear estimators, i.e., the terms $f(Z)$ in $(\mathbb{P}_{n}-\mathbb{E})\{f(Z)\}$). By combining one-sided confidence intervals for the upper and lower bounds on the ATE, one can conduct inference for the ATE itself. The following result gives the upper one-sided confidence interval for the ATE using the estimator for the upper bound $\widehat{\mathcal{U}}(\Gamma)$, while a lower one-sided interval can be constructed analogously using ^$\widehat{\mathcal{L}}(\Gamma)$.
Corollary 1.
Let $UB(\alpha)=\widehat{\mathcal{U}}(\Gamma)+\Phi^{-1}(1-\alpha)\sqrt{\frac{\mathbb{V}\{\phi_{\mathcal{U}}(Z)\}}{n}}$, where $\Phi(\cdot)$ is the cumulative distribution function of the standard normal. Suppose Assumption 1 holds, and either the setup of Theorem 1 holds and $\phi_{\mathcal{U}}$ is the influence function in (15), or the setup of Theorem 2 holds and $\phi_{\mathcal{U}}$ is the influence function in (16). Then, $\mathbb{P}\left\{\psi_{\ast}\leq UB(\alpha)\right\}\geq 1-\alpha+o(1).$
Corollary 1 demonstrates how to construct an asymptotically valid one-sided confidence interval for the ATE: $(-\infty,UB(\alpha)]$. Analogously, one can construct $LB(\alpha)$ such that $\mathbb{P}\left\{\psi_{\ast}\geq LB(\alpha)\right\}\geq 1-\alpha+o(1).$ Hence, one can construct an asymptotically valid two-sided $1-\alpha$ confidence interval for the ATE as $\left[LB(\alpha/2),UB(\alpha/2)\right]$.
Corollary 1 leverages the influence functions derived in Theorems 1 & 2 which depend on measured confounding and account for uncertainty in estimating $M$. In many cases, the limiting variance in the confidence intervals could be estimated with the sample variance of the relevant influence function $\phi_{\mathcal{U}}(Z)$. However, because the influence function could be quite complicated — particularly in Theorem 2, depending on the choice of finite-dimensional model and link function defining $\pi_{1}^{\perp}$ — one might prefer a resampling method such as the nonparametric bootstrap (Efron, 1992).
6 Illustrative Data Analysis
In this section, we illustrate our methods with a data analysis. We examine the effect of mothers’ smoking on infant birth weight. We use a calibrated sensitivity analysis to assess the robustness of causal estimates to unmeasured confounding, and compare our results to what one would conclude from a standard sensitivity analysis. In the supplementary materials, we also examine the effect of exposure to violence on attitudes towards peace in Darfur, and assess sensitivity to unmeasured confounding with both the effect differences model and the odds ratio model. Our code is available at https://github.com/alecmcclean/Calibrated-sensitivity-models.
6.1 Data
The data contains information about infant birth weights and mothers’ health in Pennsylvania between 1989 and 1991 (Almond et al., 2005; Cattaneo, 2010). It has been used to estimate the causal effect of mothers’ smoking on infant birth weight. We used a publicly available dataset consisting of 5,000 randomly subsampled observations from the original dataset (available at https://github.com/mdcattaneo/replication-C_2010_JOE). The treatment variables takes six values corresponding to ranges of cigarettes smoked daily ($0\}$ , 1 - 5 , 6 - 10 , 11 - 15 , 16 - 20 , > 20 }), which we dichotomize into an indicator for smoking. The outcome variable is the birth weight of the infant in grams. Covariates include the race and age of the mother and father, the education level each attained, the mother’s marital status and foreign born status, the county of birth, and indicators for trimester of first prenatal care visit and mother’s alcohol use.
6.2 Methods
We analyzed the data assuming the maximum leave-one-out (LOO) effect differences model (calibrated sensitivity model 1). We constructed estimators for the bounds on the ATE according to Definition 1, with $\Gamma\in\{0.5,1,\dots,5\}$. We constructed estimators for the adjusted mean difference with different covariate subsets using the npcausal package in R (Kennedy, 2023a; R Core Team, 2021). We used 5 splits and estimated the propensity score and outcome regression functions with the SuperLearner, stacking the sample average and a random forest from the ranger package with default tuning parameters (van der Laan et al., 2007; Wright and Ziegler, 2017). We constructed a pointwise confidence band for the upper bound on the ATE according to (15) and a pointwise confidence band for the lower bound analogously. As described in Section 5.3, we constructed one-sided 97.5% confidence bands for the upper and lower bounds, and took their intersection to construct a 95% confidence band for the ATE. We also conducted a standard sensitivity analysis and post hoc calibration step, where we standardized the sensitivity parameter by estimated measured confounding. The sensitivity analysis imposed the bound $\left|\psi_{\ast}-\psi\right|\leq\gamma$, while measured confounding was $M=\max_{j\in[d]}\left|\psi-\psi_{-j}\right|$.
6.3 Results
Two key point estimates in our results are the ATE estimate with all covariates and the estimated maximum LOO measured confounding. We estimated a negative and significant ATE of -272g (95% CI: $[-360g,-184g]$), which, interpreted causally, says that a mother’s smoking will lower their child’s birth weight by 272g, on average. In addition, we estimated that the parents’ educational attainment was the most impactful confounder, and that the absolute change in the adjusted mean difference without it included as a covariate was 47g (95% CI: $[0g,118g]$). There is a large degree of uncertainty in the estimate of the impact of education — the lower bound of its 95% confidence interval is zero.
Figure 1 shows the ATE estimates, estimates for bounds on the ATE, and 95% confidence bands for the ATE. The x-axis is the level of the sensitivity parameter, $\Gamma$, in the calibrated sensitivity model, and the sensitivity parameter standardized by estimated measured confounding, $\gamma/\widehat{M}$, in the standard sensitivity model. The y-axis is at the level of the causal effect. The horizontal dashed line is the ATE estimate, and the horizontal dotted line is at zero. The dot-dash lines are the estimates of the upper and lower bounds on the ATE. The point estimates for these bounds are the same in the calibrated sensitivity model and sensitivity model. However, the confidence intervals for the bounds differ. The confidence intervals for the calibrated sensitivity model are solid, while the confidence intervals for the sensitivity model are long dashes.
Remark 2.
The estimated bounds are the same in both analyses after standardizing the x-axis. We give further details as to why this occurs in Appendix A in the supplementary materials, but in short, for the sensitivity model $\left|\psi_{\ast}-\psi\right|\leq\gamma$, one can standardize the x-axis by estimated measured confounding, i.e., change the x-axis to $\gamma/\widehat{M}$, to point estimate the same bounds as the calibrated sensitivity analysis with the effect differences model.
The calibrated sensitivity results in Figure 1 can be interpreted in the following way. The estimated bounds on the ATE do not intersect zero for unmeasured confounding less than five times measured confounding, and the confidence interval for the ATE intersects zero when $\Gamma\approx 3$, so the statistical significance of the estimated effect would only be nullified for unmeasured confounding at least thrice as big as measured confounding. In this case, the measured confounding is the maximum change in the adjusted mean difference from leaving out one covariate. The most impactful observed confounder was the parent’s educational attainment, and it changed the adjusted mean difference by 47g.
While both the traditional and calibrated sensitivity analyses suggest that each effect is somewhat robust to confounding, they suggest it to differing degrees. Indeed, these results illustrate how not accounting for uncertainty in estimating measured confounding can lead to over- or under-estimating robustness to unmeasured confounding. For $\Gamma>2$, the calibrated sensitivity model confidence intervals always contain the standard sensitivity model confidence intervals. Therefore, the standard analysis over-estimates robustness to unmeasured confounding. However, for $\Gamma<2$, the calibrated CI is tighter on the upper bound, and thus could lead to concluding the effect estimate is more robust to unmeasured confounding (i.e., the standard analysis under-estimates robustness to unmeasured confounding).
Figure 1: The effect of mothers’ smoking on infant birth weight. This figure shows the estimate for the ATE, estimates for bounds on the ATE, and 95% confidence intervals for the ATE (y-axis) at different levels of unmeasured confounding (x-axis).
7 Discussion
In this paper, we introduced calibrated sensitivity models to estimate nonparametric bounds on causal effects while allowing for unmeasured confounding. Calibrated sensitivity models bound the error due to unmeasured confounding by a multiple of measured confounding. We demonstrated that calibrated sensitivity analyses solve several deficiencies of standard sensitivity and post hoc calibration analyses, including (1) the calibrated sensitivity parameter has an intuitive unit-less definition as the ratio of unmeasured over measured confounding, unlike standard sensitivity parameters; (2) uncertainty in the estimate for measured confounding can be incorporated, which prior methods have failed to do; and (3) the choice of measured confounding, which is often unjustified in standard post hoc calibration, might be better justified in practice with calibrated sensitivity models, because the metric for measured confounding is an assumption within the calibrated sensitivity model itself, rather than a choice that’s made after the sensitivity analysis is conducted. Moreover, we argued that new quantifications of unmeasured confounding, which may have advantages over standard methods (e.g., more flexible, simpler), can be bounded in an intuitive manner in a calibrated sensitivity framework. Dissemination of methods with these new models could increase the adoption of sensitivity analyses in practice.
We provided intuition for calibrated sensitivity models through several examples. Then, we identified bounds on the average treatment effect and derived methods for estimation and inference for those bounds which account for uncertainty in estimated measured confounding. We established when the estimators are $\sqrt{n}$-consistent and asymptotically normal under doubly robust style conditions on their nuisance function estimators. Finally, we conducted a data analysis to illustrate these methods, examining the effect of mothers’ smoking on infant birth weight. By correctly incorporating uncertainty due to estimating measured confounding, we observe that causal effects can less robust or more robust to unmeasured confounding than what would have been shown with standard approaches.
There are several natural directions for future work. First, other choices of unmeasured and measured confounding are possible within the partial identification framework we considered. We provide theoretical results for three different choices, and often each unique choice will require its own theoretical analysis. Therefore, an interesting avenue for future work is establishing further theoretical and inferential results for other partial identification calibrated sensitivity models. Second, our results focus on confidence intervals for the bounds of the average treatment effect, and thus a natural extension would be to examine other functions of these intervals and bounds, such as the minimum value of the sensitivity parameter that would reverse the conclusions of the analysis. Finally, two promising avenues include developing calibrated sensitivity models that point identify causal effects, and extending these methods to conduct calibrated sensitivity analyses for violations of other assumptions, such as covariate shift.
Acknowledgments
The authors thanks participants in the Carnegie Mellon University causal reading group, CMStatistics 2023, and the 2024 American Causal Inference Conference for helpful comments and feedback.
Supplementary Materials
These supplementary materials are split into seven sections:
  1. Appendix A
    discusses an invariance property of the bounds and provides an illustrative analysis demonstrating when standard sensitivity analyses with post hoc calibration over- or under-estimate robustness to unmeasured confounding.
  2. Appendix B
    provides estimation and inference results with the outcome model (calibrated sensitivity model 3), which complement those for the effect differences and odds ratio models in Section 6.
  3. Appendix C
    describes two data analyses we conducted with the effect differences and odds ratio models, assessing the effect of exposure to violence on attitudes towards peace in Darfur, and the sensitivity of that causal effect to unmeasured confounding.
  4. Appendix D
    provides results for Section 4, Proposition 1 and Lemma 1. Lemma 1 is proven by stating and proving a more detailed result, Lemma 2.
  5. Appendix E
    provides the relevant efficient influence functions for estimating measured confounding and the bounds, deriving new efficient influence functions for the outcome model.
  6. Appendix F
    provides results for Section 5 and Appendix B; in particular, for the convergence results in Theorems 1-3.
  7. Appendix G
    provides a technical result for estimating the maximum of a finite number of functionals, which is used in the proofs in Appendix F.
Appendix A Invariance of the bounds and intuition for when standard approaches under- or over-estimate robustness to unmeasured confounding
This section builds further intuition for understanding the behavior of bounds and estimators for them with calibrated sensitivity models.
A.1 Properties of the bounds
We begin with a simple observation, that the bounds in a sensitivity model and calibrated sensitivity model will be the same when $\gamma=\Gamma M$.
Proposition 2.
(Invariance) Suppose there is a sensitivity model that imposes a one-dimensional bound of the form $U\leq\gamma\in(0,\infty)$, where $U$ is some quantification of unmeasured confounding, which implies a deterministic upper bound on the causal effect of interest, $u(\gamma):(0,\infty)\to\mathbb{R}$. Moreover, suppose there is a related calibrated sensitivity model of the form $U\leq\Gamma M$, where $U$ is the same as before, $\Gamma\in(0,\infty)$, and $0<M<\infty$ (as in Assumption 3) is some quantification of measured confounding. The calibrated sensitivity model then implies a deterministic bound $\mathcal{U}(\Gamma)=u(\Gamma M)$ and, therefore,
$\Gamma M=\gamma\implies\mathcal{U}(\Gamma)=u(\gamma).$ (17)
The relationship in (17) is true under the weak condition that the bound $u(\gamma)$ is a deterministic function, so it maps the same input to the same output. It also requires that measured confounding is non-zero and bounded, so that the calibrated sensitivity bound is non-zero and non-infinite. This result holds for the effect differences and odds ratio models. For the outcome model, which implicitly imposes two bounds — for $a\in\{0,1\}$ — the result in (17) does not hold, but a similar two-dimensional property applies.
The observation in Proposition 2 is important, despite its simplicity, because it illustrates how calibrated sensitivity analyses can be compared with standard methods. Standard approaches compare $\gamma$ values to estimates $\widehat{M}$ in terms of their relative magnitude. Thus, standard approaches implicitly consider the ratio $\gamma/M$, and (17) confirms that if $\frac{\gamma}{M}=\Gamma$ then $\left\{\frac{\gamma}{M},u(\gamma)\right\}=\{\Gamma,\mathcal{U}(\Gamma)\}$. In other words, standard sensitivity models with post hoc calibration identify the same tuple as calibrated sensitivity models. Therefore, we can compare uncertainty quantification in each approach by dividing $\gamma$ by $\widehat{M}$, and comparing the confidence intervals for the second elements in $\left\{\gamma/\widehat{M},\widehat{u}(\gamma)\right\}$ and $\{\Gamma,\widehat{\mathcal{U}}(\Gamma)\}$. This comparison enables understanding when standard methods over- or under-estimate robustness to unmeasured confounding.
A.2 Understanding the change in robustness to unmeasured confounding
For simplicity, we focus on the effect differences model, although the conclusions generalize immediately to other models with bounds that are linear in the sensitivity parameter, such as the outcome model, and may generalize to other models.
With the effect differences sensitivity model, the upper bound takes the form $u(\gamma)=\psi+\gamma$, where $\psi$ is some functional (in this paper, the adjusted mean difference) that depends on the data. With the calibrated sensitivity model, the upper bound is $\mathcal{U}(\Gamma)=\psi+\Gamma M.$ Inspired by the invariance property in Proposition 2 in Section 4, we compare confidence intervals of stylized estimators that point estimate the same tuple. Suppose $\gamma=\Gamma M$. We compare confidence intervals for the second element in the following tuples:
  1. 1.
    Sensitivity model: $\left\{\gamma/\widehat{M},\widehat{u}(\gamma)\right\}$
  2. 2.
    Calibrated sensitivity model: $\left\{\Gamma,\widehat{\mathcal{U}}(\Gamma)\right\}$
Assuming appropriately accurate estimators, both approaches estimate the same tuple because $\gamma=\Gamma M$. The difference arises in how confidence intervals are constructed for the second element — the calibrated sensitivity approach accounts for uncertainty in estimating $\widehat{M}$, while the standard post hoc calibration approach does not.
Suppose access to estimators such that Wald-type confidence intervals can be constructed using the variance of each estimator:
$\mathbb{V}\left\{\widehat{u}(\gamma)\right\}$ $=\mathbb{V}(\widehat{\psi})\text{ and }$ (18)
$\mathbb{V}\left\{\widehat{\mathcal{U}}(\Gamma)\right\}$ $=\mathbb{V}(\widehat{\psi}+\Gamma\widehat{M})=\mathbb{V}\left(\widehat{\psi}+\gamma\frac{\widehat{M}}{M}\right)\equiv\mathbb{V}\left\{\widehat{u}(\gamma)+\gamma\frac{\widehat{M}}{M}\right\}$ (19)
We can compare confidence intervals by comparing the sizes of the variances in (18) and (19). If the variance of $\widehat{u}(\gamma)$ is larger than the variance of $\widehat{\mathcal{U}}(\Gamma)$, it implies that standard methods are under-estimating robustness to unmeasured confounding, in the sense that a seemingly smaller amount of unmeasured confounding would be needed to nullify a significant causal effect compared to what one would conclude from a calibrated sensitivity analysis that incorporates uncertainty in $\widehat{M}$. Similarly, if the variance of $\widehat{u}(\gamma)$ is smaller, it implies that standard methods are over-estimating robustness to unmeasured confounding. Notice that
$\frac{\mathbb{V}\{\widehat{\mathcal{U}}(\Gamma)\}}{\mathbb{V}\{\widehat{u}(\gamma)\}}=1+\frac{\mathbb{V}(\widehat{M})/M^{2}}{\mathbb{V}\{\widehat{u}(\gamma)\}/\gamma^{2}}+2\rho\left(\widehat{u}(\gamma),\widehat{M}\right)\sqrt{\frac{\mathbb{V}(\widehat{M})/M^{2}}{\mathbb{V}\{\widehat{u}(\gamma)\}/\gamma^{2}}},$ (20)
where $\rho(X,Y)=\frac{\text{cov}(X,Y)}{\sqrt{\mathbb{V}(X)\mathbb{V}(Y)}}$ is the correlation between $X$ and $Y$. The second summand on the right-hand side, $\frac{\mathbb{V}(\widehat{M})/M^{2}}{\mathbb{V}\{\widehat{u}(\gamma)\}/\gamma^{2}}$, is, roughly, the relative standardized efficiency of the estimators $\widehat{M}$ and $\widehat{u}(\gamma)$, while the third term is two times the product of the correlation between the estimators, $\rho\{\widehat{u}(\gamma),\widehat{M}\}$, and the square root of their relative standardized efficiency. Rearranging (20), we can characterize when standard methods over- or under-estimate robustness to unmeasured confounding by the following inequalities:
$\frac{1}{2}\sqrt{\frac{\mathbb{V}(\widehat{M})/M^{2}}{\mathbb{V}\{\widehat{u}(\gamma)\}/\gamma^{2}}}<-\rho\{\widehat{u}(\gamma),\widehat{M}\}\implies\mathbb{V}\{\widehat{\mathcal{U}}(\Gamma)\}<\mathbb{V}\{\widehat{u}(\gamma)\}$
$\frac{1}{2}\sqrt{\frac{\mathbb{V}(\widehat{M})/M^{2}}{\mathbb{V}\{\widehat{u}(\gamma)\}/\gamma^{2}}}>-\rho\{\widehat{u}(\gamma),\widehat{M}\}\implies\mathbb{V}\{\widehat{\mathcal{U}}(\Gamma)\}>\mathbb{V}\{\widehat{u}(\gamma)\}$ (21)
Figure 2: Visualization of (21), showing when sensitivity analyses with post hoc calibration over- or under-estimate robustness to unmeasured confounding, compared to calibrated sensitivity models.
Figure 2 helps build further intuition for when standard approaches over- or under-estimate robustness to unmeasured confounding. In the red area in the top-right of Figure 2, standard approaches over-estimate robustness to unmeasured confounding. In the blue area in the bottom-left, standard approaches under-estimate robustness to unmeasured confounding. Along the white diagonal the approaches yield the same conclusions. This plot does not characterize the whole space because it is red for $>2$ on the x-axis and $>0$ on the y-axis (i.e., for most of the space, standard approaches over-estimate robustness to unmeasured confounding).
This analysis gives some intuition for when standard methods might over- or under-estimate robustness to unmeasured confounding by not appropriately accounting for uncertainty in estimates of measured confounding. Section 6 in the main body of the paper, and Appendix C, we compare and contrast calibrated sensitivity analyses with standard sensitivity analyses using several data analyses. There we find that, indeed, results can be more or less robust to unmeasured confounding when one properly accounts for uncertainty in estimating measured confounding.
Appendix B Estimation and inference with the outcome model
Here, we consider estimating the upper bound on the ATE in the outcome model (calibrated sensitivity model 3). These are complementary results to those for the effect differences and odds ratio models in Section 5.
Definition 4.
Let $\mathcal{U}(\Gamma)$ be as in (11). Construct an estimator for the upper bound as
$\widehat{\mathcal{U}}(\Gamma):=\mathbb{P}_{n}\{\widehat{\phi}(Z)\}+\Gamma\sum_{a\in\{0,1\}}\sqrt{\mathbb{P}_{n}\{\widehat{\xi}_{1-a}(X)\}}\sqrt{\frac{1}{|\mathcal{S}|}\sum_{S\in\mathcal{S}}\mathbb{P}_{n}\{\widehat{\lambda}_{a}(Z;S)\}},$ (22)
where
  • where $\phi$ is the EIF of the adjusted mean difference, defined in (14),
  • $\xi_{a}$ is the EIF of $\lVert\pi_{a}(X)\rVert_{2}^{2}$, defined in (46) in the supplementary materials,
  • $\lambda_{a}(Z;S)$ is the EIF of $\lVert\mu_{a}(X_{-S})-\mathbb{E}\big{\{}\mu_{a}(X)\mid A=1-a,X_{-S}\}\rVert_{2}^{2}$, defined in (47) in the supplementary materials, and
  • the estimated nuisance functions constituting $\widehat{\phi},\widehat{\xi}$, and $\widehat{\lambda}$ are constructed on a separate sample.
The estimator in (22) has the same structure as the estimator in (13) for the effect differences model — the first term comes from estimating the adjusted mean difference while the other terms come from estimating the bound on $\left|\psi_{\ast}-\psi\right|$ implied by the model, i.e., the second summand in the partial identification result in (11) in Proposition 1. The form of the other terms follows from estimating the two functionals in the bound, $\lVert\pi_{1-a}(X)\rVert_{2}$ and $\sqrt{\frac{1}{|\mathcal{S}|}\sum_{S\in\mathcal{S}}\left\lVert\mu_{a}(X_{-S})-\mathbb{E}\big{\{}\mu_{a}(X)\mid A=1-a,X_{-S}\}\right\rVert_{2}^{2}}$.
A crucial step in constructing the estimator for measured confounding is to construct an estimator for $\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-S}\}$, which appears in the definition of the bound in (11). To construct an estimator for $\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-S}\}$, one could use a two-stage estimator where the EIF of $\mathbb{E}\{\mu_{a}(X)\}$ is estimated and regressed against $A=1-a,X_{-S}$ as a pseudo-outcome in a second-stage regression [Kennedy, 2023b, McClean et al., 2024, Fisher and Fisher, 2023, Semenova and Chernozhukov, 2021, Díaz et al., 2018]. Therefore, in the next result, we let $\widehat{\mathbb{E}}_{n}\{\widehat{\mu}_{a}(X)\mid A=1-a,X_{-S}\}$ denote a generic regression estimator that regresses estimated EIF values of $\mathbb{E}\{\mu_{a}(X)\}$ against $1-a$ and $X_{-S}$. Linear smoothers as second-stage regressions have been studied in this context, and could achieve $n^{-1/4}$ rates when $\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-S}\}$ is appropriately smooth as a function of $X_{-S}$ [Kennedy, 2023b]. Some analyses of two-stage estimators require further sample splitting, so that the nuisance functions constituting the EIF of $\mathbb{E}\{\mu_{a}(X)\}$ are estimated on a separate sample from the second-stage regression. We will ignore that complication here, for ease of exposition.
Theorem 3.
(Average LSO outcome model) Let $\mathcal{U}(\Gamma)$ and $\widehat{\mathcal{U}}(\Gamma)$ be as in Definition 4. Suppose Assumptions 2-3 hold, and
  1. 1.
    for $S\in\mathcal{S}\cup\emptyset$, $\widehat{\pi}_{1}(X_{-S})$ is bounded away from zero and one,
  2. 2.
    $\widehat{\phi}$ is consistent in the sense that $\lVert\widehat{\phi}(Z)-\phi(Z)\rVert_{2}\stackrel{{\scriptstyle p}}{{\rightarrow}}0$,
  3. 3.
    for $a\in\{0,1\}$, $\widehat{\xi}_{a}$ is consistent in the sense that $\lVert\widehat{\xi}_{a}(X)-\xi_{a}(X)\rVert_{2}\stackrel{{\scriptstyle p}}{{\rightarrow}}0$,
  4. 4.
    for $a\in\{0,1\}$ and $S\in\mathcal{S}$, $\widehat{\lambda}_{a}(Z;S)$ is consistent in the sense that $\lVert\widehat{\lambda}_{a}(Z;S)-\lambda_{a}(Z;S)\rVert_{2}\stackrel{{\scriptstyle p}}{{\rightarrow}}0$,
  5. 5.
    for $a\in\{0,1\}$, $\lVert\widehat{\pi}_{1}(X)-\pi_{1}(X)\rVert_{2}\lVert\mu_{a}(X)-\mu_{a}(X)\rVert_{2}=o_{\mathbb{P}}(n^{-1/2})$,
  6. 6.
    $\lVert\widehat{\pi}_{1}(X)-\pi_{1}(X)\rVert_{2}=o_{\mathbb{P}}(n^{-1/4})$, and
  7. 7.
    for $a\in\{0,1\}$ and $S\in\mathcal{S}$,
    $o_{\mathbb{P}}(n^{-1/2})$ $=\left(\left\lVert\widehat{\mu}_{a}(X_{-S})-\mu_{a}(X_{-S})\right\rVert_{2}+\left\lVert\widehat{\mathbb{E}}_{n}\{\widehat{\mu}_{a}(X)\mid A=1-a,X_{-S}\}-\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-S}\}\right\rVert_{2}\right)^{2}$
    $+\lVert\widehat{\pi}_{1-a}(X_{-S})-\pi_{1-a}(X_{-S})\rVert_{2}\lVert\widehat{\mu}_{a}(X_{-S})-\mu_{a}(X_{-S})\rVert_{2}$
    $+\lVert\widehat{\pi}_{1-a}(X_{-S})-\pi_{1-a}(X_{-S})\rVert_{2}\left\lVert\mathbb{E}\{\widehat{\mu}_{a}(X)\mid A=1-a,X_{-S}\}-\widehat{\mathbb{E}}_{n}\{\widehat{\mu}_{a}(X)\mid A=1-a,X_{-S}\}\right\rVert_{2}$
    $+\lVert\widehat{\mu}_{a}(X)-\mu_{a}(X)\rVert_{2}\left(\left\lVert\widehat{\pi}_{1-a}(X)-\pi_{1-a}(X)\right\rVert_{2}+\left\lVert\widehat{\pi}_{1-a}(X_{-S})-\pi_{1-a}(X_{-S})\right\rVert_{2}\right).$ (23)
Then,
$\widehat{\mathcal{U}}(\Gamma)-\mathcal{U}(\Gamma)=(\mathbb{P}_{n}-\mathbb{E})\Bigg{(}\phi(Z)+\Gamma\Bigg{[}\sum_{a\in\{0,1\}}$ $\left\lVert\pi_{1-a}(X)\right\rVert_{2}\left\{\frac{1}{|\mathcal{S}|}\sum_{S\in\mathcal{S}}\frac{\lambda_{a}(Z;S)}{2M_{a}}\right\}$
$+M_{a}\left\{\frac{\xi_{1-a}(X)}{2\lVert\pi_{1-a}(X)\rVert_{2}}\right\}\Bigg{]}\Bigg{)}+o_{\mathbb{P}}(n^{-1/2}).$ (24)
where
$M_{a}=\sqrt{\frac{1}{|\mathcal{S}|}\sum_{S\in\mathcal{S}}\lVert\mu_{a}(X_{-S})-\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-S}\}\rVert_{2}^{2}}.$
Theorem 3 shows that the error of the estimator for the upper bound in the outcome model behaves like a centered sample average plus asymptotically negligible error under doubly robust conditions. Here, we give some intuition for the conditions. Condition 1 ensures that the bias of $\widehat{\mathcal{U}}(\Gamma)$ is bounded. Conditions 2, 3, and 4 are weak consistency conditions for controlling the empirical process terms that arise in estimating the three components of the bound: $\psi$, $\lVert\pi_{a}(X)\rVert_{2}$, and $\sqrt{\frac{1}{|\mathcal{S}|}\sum_{S\in\mathcal{S}}\lVert\mu_{a}(X_{-S})-\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-S}\}\rVert_{2}^{2}}$. These conditions would hold under weak nonparametric assumptions. Condition 5 is the usual doubly robust condition for controlling the conditional bias for estimating the adjusted mean difference $\psi$. Conditions 6 and 7 are stronger conditions, for controlling the other terms in the bias of $\widehat{\mathcal{U}}(\Gamma)$. Condition 6 controls the conditional bias of the estimator for $\lVert\pi_{a}(X)\rVert_{2}$, while condition 7 controls the conditional bias for estimating $\sqrt{\frac{1}{|\mathcal{S}|}\sum_{S\in\mathcal{S}}\lVert\mu_{a}(X_{-S})-\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-S}\}\rVert_{2}^{2}}$. Both are doubly-robust style conditions — they consist of products of errors, and thus a sufficient condition for each to hold is that each nuisance function is estimated at only an $n^{-1/4}$ rate.
Remark 3.
When the measured confounding is an average, as in the average LSO outcome model, $\sqrt{n}$ convergence guarantees on the nuisance function estimators are required across all covariate sets under consideration (in condition 7). This is necessary so that none of the biases from estimating each sub-quantification of measured confounding dominate asymptotically. This is different from the maximum leave-one-out confounding, where $\sqrt{n}$ convergence guarantees are only required for the sub-quantification of measured confounding corresponding to the maximum (e.g., condition 3 in Theorem 1). In this sense, it is easier to estimate maximum measured confounding than average measured confounding, conditional on there being a unique maximum. However, the estimation procedures are the same whether measured confounding is a maximum or an average, in the sense that one would still aim to estimate the nuisance functions as accurately as possible with each covariate subset even when measured confounding is a maximum. This is because, a priori, one cannot know which covariate subset corresponds to the true maximum, and therefore it is still necessary to estimate all nuisance functions well in both cases.
Appendix C Additional data analyses
Here, we provide two additional data analyses, assessing the effect of exposure to violence on attitudes towards peace in Darfur.
C.1 Data: peace attitudes in Darfur
The dataset contains information about the attitudes towards peace of Darfurian villagers exposed to violence. In 2003 and 2004, the Darfurian government led a series of violent attacks against civilians, killing an estimated two hundred thousand people. Due to the quasi-randomness of the attacks, with this dataset one can attempt to answer the question of whether being directly injured or maimed in such attacks made people more or less likely to accept peace [Hazlett, 2020]. We use an example version of this dataset, which is publicly available in the sensemakr package in R [Cinelli et al., 2020].
The data contains information on 1,276 Darfurian villagers. The treatment is a binary variable indicating whether the individual was physically injured during an attack, and the outcome is an index (continuous from 0 to 1) measuring pro-peace attitudes (with higher being more pro-peace). The covariates include the original village of the respondent (we grouped every village with under 10 respondents into “Other”), gender, age (in years), whether they were a herder or a farmer, whether they voted in an earlier election before the conflict, and the size of their household.
C.2 Methods
We estimated bounds on the ATE using the maximum leave-one-out effect differences model (calibrated sensitivity model 1) and the maximum leave-one-out odds ratio model (calibrated sensitivity model 2). For the effect differences model, we used the same methods as in Section 6.2 of the body of the paper. Specifically, we constructed estimators for the bounds on the ATE according to Definition 1, with $\Gamma\in\{0.5,1,\dots,5\}$. We constructed estimators for the adjusted mean difference with different covariate subsets using the npcausal package in R [Kennedy, 2023a, R Core Team, 2021]. We used 5 splits and estimated the propensity score and outcome regression functions with the SuperLearner, stacking the sample average and a random forest from the ranger package with default tuning parameters [van der Laan et al., 2007, Wright and Ziegler, 2017]. We constructed a pointwise confidence band for the upper bound on the ATE according to (15) and a pointwise confidence band for the lower bound analogously. As described in Section 5.3, we constructed one-sided 97.5% confidence bands for the upper and lower bounds, and took their intersection to construct a 95% confidence interval for the ATE. We also conducted a standard sensitivity analysis and post hoc calibration step, where we standardized the sensitivity parameter by estimated measured confounding.
In the odds ratio model, we constructed estimators for the bounds on the ATE according to Definitions 2 and 3, with $\exp(\Gamma)\in\{1.25,1.5,1.75,2\}$. We estimated the best projection of measured confounding using a logistic regression with no interactions (see Definition 2). We also assumed that the observed maxima and minima in the covariate data were the true bounds of the covariate support, which allowed us to extend Theorem 2 to a known and bounded support — in this case, we estimated measured confounding as the maximum absolute value across covariates of each covariate’s estimated coefficient multiplied by its range. We then estimated the upper bound on the ATE using the algorithm in Yadlowsky et al. [2022], with sieve estimators for the nuisance functions. We constructed one-sided 97.5% confidence bands for the upper and lower bounds and a 95% confidence band for the ATE using Wald-type confidence intervals. We estimated the limiting variance with the nonparametric bootstrap, with $100$ resampling iterations of size $1,000$.
In both analyses, we also conducted a standard sensitivity analysis and post hoc calibration step, where we standardized the sensitivity parameter by estimated measured confounding.
C.3 Results
Here, we report the results from each analysis. The point estimate and confidence interval for the ATE assuming no unmeasured confounding are the same. We estimated a positive and significant ATE of $0.06$ (95%: CI $[0.02,0.1]$) which, interpreted causally, says that exposure to violence increased people’s preference of peace by 0.06 on average (0.06 is a unit-less value — the outcome is an index continuous on $[0,1]$).
With the effect differences model, we estimated that the respondent’s original village was the most impactful confounder, and that the absolute change in the adjusted mean difference without it included as a covariate was $0.01$ with 95% CI $[0.002,0.02]$. Meanwhile, with the odds ratio model, we again estimated that the respondent’s original village was the most impactful confounder. The maximum estimated confounding is $3.34$ (95% CI: $[0,20]$) (this is a Wald-type confidence interval truncated below at 00 where the variance was calculated across $100$ bootstrap resamples). Note also that the respondent’s village was the most impactful measured confounder in 99 out of the 100 bootstrap resamples.
Figure 3 shows the ATE estimates, estimates for bounds on the ATE, and 95% confidence bands for the ATE — Figure 3(a) is for the effect differences model while Figure 3(b) is for the odds ratio model. The x-axis is the level of the sensitivity parameter, $\Gamma$, in the calibrated sensitivity model, and the sensitivity parameter standardized by estimated measured confounding, $\gamma/\widehat{M}$, in the standard sensitivity model. The y-axis is at the level of the causal effect. The horizontal dashed lines are the ATE estimates, and the horizontal dotted lines are at zero. The dot-dash lines are the estimates of the upper and lower bounds on the ATE. By the invariance property discussed in Proposition 2 in Section 4, the point estimates for these bounds are the same in the calibrated sensitivity model and sensitivity model. However, the confidence intervals for the bounds differ. The confidence intervals for the calibrated sensitivity model are solid, while the confidence intervals for the sensitivity model are long dashes.
The calibrated sensitivity results in Figure 3(a) can be interpreted in the following way. The estimated bounds on the ATE do not intersect zero for unmeasured confounding less than five times measured confounding, and the confidence interval for the ATE intersects zero when $\Gamma\approx 2$, so the statistical significance of the estimated effect would only be nullified for unmeasured confounding at least twice as big as measured confounding. In this case, the measured confounding is the maximum change in the adjusted mean difference from leaving out one covariate. The most impactful observed confounder was the respondent’s home village, and it changed the adjusted mean difference by 0.01. The results in Figure 3(b) can be interpreted in a similar way.
The relationship between the calibrated sensitivity results and standard sensitivity results are similar in each analysis, in the sense that incorporating uncertainty due to estimating measured confounding demonstrates that the causal effect estimate is less robust to unmeasured confounding (i.e., the solid confidence intervals encompass the dashed confidence intervals). However, with both the standard and calibrated sensitivity models, the odds ratio model suggests the estimates are far less robust to unmeasured confounding than what is suggested by the effect differences model.
(a) Effect differences model.
(b) Odds ratio model.
Figure 3: The effect of exposure to violence on preference for peace in Darfur. These figures show estimates for the ATE, estimates for bounds on the ATE, and 95% confidence intervals for the ATE (y-axis) at different levels of unmeasured confounding (x-axis).
C.4 Discussion: the odds ratio model
Here, we further discuss the results from the odds ratio model. Without standardizing by measured confounding, the confidence interval in the standard sensitivity model intersects the x-axis for $\gamma\approx 0.11$ (multiplying the point where the long dashed lines in Figure 3(b) intersect the x-axis by the estimate for measured confounding $\widehat{M}$). On the exponential scale, which this model is usually expressed in, that corresponds to a sensitivity parameter of roughly $1.12$. Arguably, this indicates much less robustness to unmeasured confounding than is observed in the the effect differences analysis. The lack of robustness to unmeasured confounding may be an artifact of the estimation process — for computational simplicity, we manually constructed sieve estimators for the nuisance functions $\nu_{a}^{\pm}$ and $\theta_{a}^{\pm}$ using a linear model with no interactions. This may be a poor approximation of the true functions, and better estimators could lead to different results.
Moreover, when measured confounding is incorporated in the calibrated sensitivity model, the results become even less robust to unmeasured confounding — the calibrated confidence intervals intersect the x-axis for $\Gamma\approx 0.025$, suggesting unmeasured confounding which is only a small multiple of measured confounding would be enough to overturn the conclusions of the study. Part of the reason this occurs is because the maximum LOO confounding is high and has high variance. The maximum estimated confounding is $3.34$, with a huge confidence interval of $[0,20]$ (this is a Wald-type confidence interval truncated below at 00 where the variance was calculated across $100$ bootstrap resamples). The estimated measured confounding may be high because the best projection estimator is a poor approximation of the true propensity score. The high variance across bootstrap resamples reinforces this and suggests this estimator is unstable. More generally, understanding why these results indicate such different conclusions from the effect differences analysis is a topic of future work.
Appendix D Results in Section 4
Proposition 1
Proof:
The proof follows for the effect differences model directly from the definition of the model, while for the odds ratio model it follows by Yadlowsky et al. [2022] Lemma 2.1.
For the outcome model,
$\mathbb{E}(Y^{1}-Y^{0})$ $=\mathbb{E}\left\{\mathbb{E}(Y^{1}-Y^{0}\mid X)\right\}$
$=\mathbb{E}\left\{\mathbb{E}(Y^{1}-Y^{0}\mid A=1,X)\pi_{1}(X)+\mathbb{E}(Y^{1}-Y^{0}\mid A=0,X)\pi_{0}(X)\right\}$
$=\mathbb{E}\{\mathbb{E}(Y^{1}\mid A=1,X)-\mathbb{E}(Y^{0}\mid A=0,X)\}$
$\hskip 14.45377pt-\mathbb{E}\left\{\mathbb{E}(Y^{1}\mid A=1,X)\pi_{0}(X)+\mathbb{E}(Y^{0}\mid A=1,X)\pi_{1}(X)\right\}$
$\hskip 14.45377pt+\mathbb{E}\left\{\mathbb{E}(Y^{1}\mid A=0,X)\pi_{0}(X)+\mathbb{E}(Y^{0}\mid A=0,X)\pi_{1}(X)\right\}$
$=\psi+\sum_{a\in\{0,1\}}(2a-1)\mathbb{E}\left[\left\{\mathbb{E}(Y^{a}\mid A=1-a,X)-\mu_{a}(X)\right\}\pi_{1-a}(X)\right].$
Therefore, by Hölder’s inequality,
$\left|\psi^{\ast}-\psi\right|\leq\sum_{a\in\{0,1\}}\Big{\lVert}\mathbb{E}(Y^{a}\mid A=1-a,X)-\mu_{a}(X)\Big{\rVert}_{p}\Big{\lVert}\pi_{1-a}(X)\Big{\rVert}_{r}$
where $r=1$ if $p=\infty$ and $r=\frac{p}{p-1}$ otherwise. Under calibrated sensitivity model 3, the result follows by setting $p=r=2$ and imposing the model. $\square$
Proposition 2
Proof:
The calibrated sensitivity model imposes the bound $U\leq\Gamma M$. Because $\Gamma\in(0,\infty)$ and $0<M<\infty$ by assumption, $\Gamma M\in(0,\infty)$. Therefore, this model implies an upper bound on the causal effect of interest: $u(\Gamma M)$ (substitute $\gamma^{\prime}=\Gamma M$ and use the assumption/construction of the sensitivity model). Because $u(\cdot)$ is a deterministic function, it maps the same inputs to the same outputs, and the result follows. $\square$
D.1 Differentiability with the odds ratio model
Here, we provide the details on the derivatives considered implicitly in Lemma 1. For analyzing the odds ratio model, we require several mild continuity and boundedness assumptions. For this purpose, we define one more nuisance function and a loss function:
$\omega_{\theta}(Y;t)$ $:=\big{(}Y-\theta\big{)}\mathbbold{1}\big{(}Y>\theta\big{)}+t\big{(}Y-\theta\big{)}\mathbbold{1}\big{(}Y<\theta\big{)},$ (25)
$\ell_{t}^{-}(\theta,y)$ $:=\frac{1}{2}\Big{[}\big{\{}(y-\theta)\mathbbold{1}(y>\theta)\big{\}}^{2}+t\big{\{}(y-\theta)\mathbbold{1}(y<\theta)\}^{2}\Big{]}\text{, and}$ (26)
$\ell_{t}^{+}(\theta,y)$ $=\ell_{1/t}^{-}(\theta,y).$ (27)
Next, we give the continuity and boundedness conditions required (see Lemmas 2.2 and 2.3 and Assumption A.1 in Yadlowsky et al. [2022]).
Assumption 7.
Smooth and bounded data generating process:
  1. 1.
    $f(x)$, the covariate density, $\pi_{1}(x)$, and $\theta_{a}^{-}(x;t)$ (defined in (9)) and $\theta_{a}^{+}(x;t)$ for $a\in\{0,1\}$ are continuous in $x$,
  2. 2.
    for $a\in\{0,1\}$, $p(y\mid A=a,x)$, the conditional density of $Y$ given $A=a$ and $X=x$ is continuous in $y$ and $x$, upper bounded, and has bounded support,
  3. 3.
    $Y$ is bounded,
  4. 4.
    for $a\in\{0,1\}$ and all $x\in\mathbb{R}^{d}$, $\theta_{a}^{-}(x;t)$ and $\theta_{a}^{+}(x;t)$ are bounded,
  5. 5.
    for $a\in\{0,1\}$ and bounded $t$, $\theta_{a}^{-}(x;t)$ and $\theta_{a}^{+}(x;t)$ are absolutely integrable in $x$, and
  6. 6.
    for $a\in\{0,1\}$, $(r,x)\mapsto\mathbb{E}\left\{\ell_{\exp(\Gamma M)}^{-}\big{(}r,Y\big{)}\mid A=a,x\right\}$ and $(r,x)\mapsto\mathbb{E}\left\{\ell_{\exp(\Gamma M)}^{+}\big{(}r,Y\big{)}\mid A=a,x\right\}$ are continuous on ×$\mathbb{R}\times\mathbb{R}^{d}$, and $\mathbb{E}\left(\ell_{\exp(\Gamma M)}^{-}\Big{[}\theta_{a}^{-}\{X;\exp(\Gamma M)\},Y\Big{]}\mid A=a\right)<\infty$ and
    $\mathbb{E}\left(\ell_{\exp(\Gamma M)}^{+}\Big{[}\theta_{a}^{+}\{X;\exp(\Gamma M)\},Y\Big{]}\mid A=a\right)<\infty$.
This assumption imposes mild continuity and boundedness conditions. Condition 1, the continuity of $p(y\mid A=a,x)$, the boundedness of $Y$, and the absolute integrability of $\theta_{a}^{\pm}$ are beyond what is assumed in Yadlowsky et al. [2022]. Condition 2 is contained in Assumption A.1 of that paper, while condition 4 is in Lemma 2.2 and condition 6 is in Lemma 2.3. With this assumption in hand, we can now state and prove the following result — Lemma 2 — which implies Lemma 1.
Lemma 2.
(Differentiable bounds in the odds ratio model) Under the setup of Lemma 1,
$\frac{\partial}{\partial M}\mathcal{L}(\Gamma)$ $=-\Gamma\exp(\Gamma M)\mathbb{E}\left\{\frac{\pi_{0}(X)f_{1}(X;\theta_{1}^{-})}{\nu_{1}^{-}(X)}\right\}-\Gamma\mathbb{E}\left\{\pi_{1}(X)\frac{\widetilde{f}_{0}(X;\theta_{0}^{+})}{\nu_{0}^{+}(X)}\right\}\text{ and }$ (28)
$\frac{\partial}{\partial M}\mathcal{U}(\Gamma)$ $=\Gamma\mathbb{E}\left\{\pi_{0}(X)\frac{\widetilde{f}_{1}(X;\theta_{1}^{+})}{\nu_{1}^{+}(X)}\right\}+\Gamma\exp(\Gamma M)\mathbb{E}\left\{\pi_{1}(X)\frac{f_{0}(X;\theta_{0}^{-})}{\nu_{0}^{-}(X)}\right\},$ (29)
where $X\equiv X$; $\theta_{a}^{\pm}\equiv\theta_{a}^{\pm}\{X;\exp(\Gamma M)\}$; $\theta_{a}^{\pm}$ and $M$ are defined in (9) and (10), respectively; and
$f_{a}(x;\theta)$ $=\mathbb{E}\left\{\left(\theta-Y\right)\mathbbold{1}\left(\theta>Y\right)\mid A=a,X=x\right\},$ (30)
$\widetilde{f}_{a}(x;\theta)$ $=\mathbb{E}\big{\{}(Y-\theta)\mathbbold{1}(Y>\theta)\mid A=a,X=x\big{\}},$ (31)
$\nu_{a}^{-}(x)\equiv\nu_{a}^{-}\{x;\theta,t=\exp(\Gamma M)\}$ $=\mathbb{P}(Y\geq\theta\mid A=a,X=x)+t\mathbb{P}(Y<\theta\mid A=a,X=x)\text{, and}$ (32)
$\nu_{a}^{+}(x)\equiv\nu_{a}^{+}\{x;\theta,t=\exp(\Gamma M)\}$ $=\mathbb{P}(Y>\theta\mid A=a,X=x)+\frac{1}{t}\mathbb{P}(Y<\theta\mid A=a,X=x).$ (33)
Moreover, the derivative of the upper bound is positive and bounded while the the derivative of the lower bounds is negative and bounded.
Proof:
The derivatives for the upper and lower bounds on $\mathbb{E}(Y^{1})$ follow from this proof and the four lemmas afterward. The other derivatives for the upper and lower bounds on $\mathbb{E}(Y^{0})$ follow the same argument, but replacing $A=1$ by $A=0$.
Note that, by assumption $0<\exp(\Gamma M)<\infty$. Hence, Lemmas 34, and 5 apply, and the result then follows by the chain rule, Assumption 2 and the dominated convergence theorem (to bring the derivative inside the expectation), and canceling terms. The boundedness of the derivative follows because $\widetilde{f}_{a},f_{a},\nu_{a}^{\pm}$ are all bounded and $M$ and $\Gamma$ are non-zero and bounded by assumption. The sign follows because $f_{a},\widetilde{f}_{a},\nu_{a}^{\pm}>0$. $\square$
Lemma 3.
Under the setup of Lemma 2, for $t^{\prime}>0$,
$\frac{\partial}{\partial t}\theta_{1}^{-}(X;t)\Big{|}_{t=t^{\prime}}=-\frac{f_{1}(X;\theta_{1}^{-})^{2}}{g_{1}\big{(}X;\theta_{1}^{-}\big{)}},$ (34)
where $X\equiv X$, $\theta_{1}^{-}:=\theta_{1}^{-}(X;t^{\prime})$ on the right-hand side, $f_{a}(x;\theta)$ is defined in (30), and
$g_{a}(x;\theta)=f_{a}(x;\theta)\mathbb{P}\left(Y>\theta\mid A=a,X=x\right)+\widetilde{f}_{a}(x;\theta)\mathbb{P}\left(Y<\theta\mid A=a,X=x\right),$ (35)
where $\widetilde{f}$ is defined in (31).
Proof:
We establish the result using the inverse function theorem, where we express $t$ as a function of $\theta$.
Under Assumption 7, Lemma 2.3 from Yadlowsky et al. [2022] applies, such that $\theta_{1}^{-}(X;t)$ is the unique minimizer (up to measure-zero sets) of $\mathbb{E}\{\ell_{t}^{-}(\theta,Y)\mid A=1,X\}$ and solves
$\mathbb{E}\{\omega_{\theta_{1}^{-}(X;t)}(Y;t)\mid A=1,X\}=0.$ (36)
In what follows let $\theta\equiv\theta_{1}^{-}(X;t)$ to further simplify notation. The moment condition implies
$t(\theta)=\frac{\mathbb{E}\big{\{}(Y-\theta)\mathbbold{1}(Y>\theta)\mid A=1,X=x\big{\}}}{\mathbb{E}\big{\{}(\theta-Y)\mathbbold{1}(\theta>Y)\mid A=1,X=x\big{\}}}.$ (37)
It is possible to divide by $\mathbb{E}\big{\{}(\theta-Y)\mathbbold{1}(\theta>Y)\mid A=1,X=x\big{\}}$ when $t>0$ because
  1. 1.
    $\mathbb{E}\big{\{}(Y-\theta)\mathbbold{1}(Y>\theta)\mid A=1,X=x\big{\}}$ and $\mathbb{E}\big{\{}(\theta-Y)\mathbbold{1}(\theta>Y)\mid A=1,X=x\big{\}}$ are both non-negative by definition, and
  2. 2.
    $\mathbb{E}\big{\{}(Y-\theta)\mathbbold{1}(Y>\theta)\mid A=1,X=x\big{\}}$ and $\mathbb{E}\big{\{}(\theta-Y)\mathbbold{1}(\theta>Y)\mid A=1,X=x\big{\}}$ cannot each be zero simultaneously because $p(y\mid A=1,x)$ is continuous and upper bounded, which implies $Y\mid A=1,X=x$ is not a point mass.
From 1 and 2, we conclude both $\mathbb{E}\big{\{}(Y-\theta)\mathbbold{1}(Y>\theta)\mid A=1,X=x\big{\}}$ and $\mathbb{E}\big{\{}(\theta-Y)\mathbbold{1}(\theta>Y)\mid A=1,X=x\big{\}}$ must be positive.
By the continuity assumption on $p(y\mid A=1,x)$ and the quotient rule and Leibniz’ integral rule (see Lemma 6 for specific steps), we can take the derivative of $t(\theta)$ wrt $\theta$, and find
$\frac{\partial}{\partial\theta}t(\theta):=-\Bigg{(}$ $\frac{\mathbb{E}\big{\{}(\theta-Y)\mathbbold{1}(\theta>Y)\mid A=1,X=x\big{\}}\mathbb{P}(Y>\theta\mid A=1,X=x)}{\Big{[}\mathbb{E}\big{\{}(\theta-Y)\mathbbold{1}(\theta>Y)\mid A=1,X=x\big{\}}\Big{]}^{2}}$
$+\frac{\mathbb{E}\big{\{}(Y-\theta)\mathbbold{1}(Y>\theta)\mid A=1,X=x\big{\}}\mathbb{P}(Y\leq\theta\mid A=1,X=x)}{\Big{[}\mathbb{E}\big{\{}(\theta-Y)\mathbbold{1}(\theta>Y)\mid A=1,X=x\big{\}}\Big{]}^{2}}\Bigg{)}$ (38)
Notice that the denominator is positive by the prior argument and steps 1 and 2 above. Meanwhile, the numerator is also positive by the same argument and steps 1 and 2 above, and because $\mathbb{P}(Y>\theta\mid A=1,X=x)\vee\mathbb{P}(Y\leq\theta\mid A=1,X=x)>0$. Moreover, both the numerator and denominator are bounded by Assumption 7.
Hence, $\frac{\partial}{\partial\theta}t(\theta)$ exists and is always negative and bounded. Therefore, the result follows by the inverse function theorem for the reciprocal of (38). Moreover, $\frac{\partial}{\partial t}\theta(t)$ is continuous because $p(y\mid A=1,x)$ is continuous and $\frac{\partial}{\partial\theta}t(\theta)$ is a composition of integrals of $p(y\mid A=1,x)$. $\square$
Lemma 4.
Under the setup of Lemma 2, then
$\nu_{1}^{-}\big{(}X;\theta_{1}^{-},t\big{)}f_{1}(X;\theta_{1}^{-})=g_{1}\big{(}X;\theta_{1}^{-},t\big{)}$ (39)
where $X\equiv X$, $\theta_{1}^{-}:=\theta_{1}^{-}(X;t)$, and $\nu_{1}^{-}(X,\theta,t)$, $f_{a}(x;\theta)$, and $g_{a}(X;\theta,t)$ are defined in (32), (30), and (35), respectively.
Proof:
Omitting arguments, on the left-hand side of (39) we have
$\nu f_{1}=f_{1}\mathbb{P}(Y\geq\theta\mid A=1,X=x)+tf_{1}\mathbb{P}(Y\leq\theta\mid A=1,X=x).$ (40)
For the second summand on the right-hand side of (40),
$tf_{1}\mathbb{P}(Y\leq\theta\mid A=1,X=x)$ $=t\mathbb{E}\left\{\left(\theta-Y\right)\mathbbold{1}\left(\theta>Y\right)\mid A=a,X=x\right\}\mathbb{P}(Y\leq\theta\mid A=1,X=x)$
$=-t\mathbb{E}\left\{\left(Y-\theta\right)\mathbbold{1}\left(\theta>Y\right)\mid A=a,X=x\right\}\mathbb{P}(Y\leq\theta\mid A=1,X=x)$
$=\mathbb{E}\left\{\left(Y-\theta\right)\mathbbold{1}\left(Y>\theta\right)\mid A=a,X=x\right\}\mathbb{P}(Y\leq\theta\mid A=1,X=x)$
where the third equality follows because $\theta$ solves the moment equation $\mathbb{E}(\omega_{\theta}(Y)\mid A=1,X=x)=0$ (see, (36)). Hence,
$\nu f_{1}$ $=f_{1}\mathbb{P}(Y\geq\theta\mid A=1,X=x)$
$+\mathbb{E}\left\{\left(Y-\theta\right)\mathbbold{1}\left(Y>\theta\right)\mid A=a,X=x\right\}\mathbb{P}(Y\leq\theta\mid A=1,X=x)$
$=g_{1}(X;\theta).$
$\square$
Lemma 5.
Under the setup of Lemma 2, for $t^{\prime}>0$,
$\frac{\partial}{\partial t}\theta_{1}^{+}(X;t)\Big{|}_{t=t^{\prime}}=\frac{\widetilde{f}_{1}(X;\theta_{1}^{+})^{2}}{g_{1}(X;\theta_{1}^{+})}=\frac{\widetilde{f}_{1}(X;\theta_{1}^{+})}{t^{\prime}\nu_{1}^{+}\big{(}X;\theta_{1}^{+},t^{\prime}\big{)}},$ (41)
where $X\equiv X$, $\theta_{1}^{+}:=\theta_{1}^{+}(X;t^{\prime})$ on the right-hand side, and $\widetilde{f}_{a}(x;\theta)$, $g_{a}(x;\theta)$ and $\nu_{a}^{+}$ are defined in (31), (35), and (33), respectively.
Proof:
By the same argument as in Yadlowsky et al. [2022] (see, Section 2.3), $\theta_{1}^{+}(X;t)$ minimizes $\mathbb{E}\{\ell_{t}^{+}(\theta,Y)\mid A=1,X\}=\mathbb{E}\{\ell_{1/t}^{-}(\theta,Y)\mid A=1,X\}$ and solves the moment condition
$\mathbb{E}\left\{\left(Y-\theta\right)\mathbbold{1}\left(Y>\theta\right)\mid A=1,X=x\right\}-\frac{1}{t}\mathbb{E}\left\{\left(\theta-Y\right)\mathbbold{1}\left(\theta>Y\right)\mid A=1,X=x\right\}=0.$
Therefore, the first result follows by the same arguments as in Lemma 3 using the inverse function theorem.
For the second result, omitting $X$ arguments and letting $\theta\equiv\theta_{1}^{+}$,
$\frac{\widetilde{f}_{1}(\theta;t^{\prime})}{\frac{\partial}{\partial y}\theta_{1}^{+}}$ $=\frac{g_{1}(\theta;t^{\prime})}{\widetilde{f}_{1}(\theta;t^{\prime})}$
$=\frac{\mathbb{E}\left\{\left(\theta-Y\right)\mathbbold{1}\left(\theta>Y\right)\mid A=1,X=x\right\}\mathbb{E}\left(Y\geq\theta\mid A=1,X=x\right)}{\mathbb{P}\left\{\left(Y-\theta\right)\mathbbold{1}\left(Y>\theta\right)\mid A=1,X=x\right\}}$
$+\frac{\mathbb{E}\left\{\left(Y-\theta\right)\mathbbold{1}\left(Y>\theta\right)\mid A=1,X=x\right\}\mathbb{P}\left(Y\leq\theta\mid A=1,X=x\right)}{\mathbb{E}\left\{\left(Y-\theta\right)\mathbbold{1}\left(Y>\theta\right)\mid A=1,X=x\right\}}$
$=\frac{\mathbb{E}\left\{\left(\theta-Y\right)\mathbbold{1}\left(\theta>Y\right)\mid A=1,X=x\right\}}{\mathbb{E}\left\{\left(Y-\theta\right)\mathbbold{1}\left(Y>\theta\right)\mid A=1,X=x\right\}}\mathbb{P}\left(Y\geq\theta\mid A=1,X=x\right)+\mathbb{P}(Y\leq\theta\mid A=1,X=x).$
Because $\theta_{1}^{+}$ solves the moment equation
$\mathbb{E}\left\{\left(Y-\theta\right)\mathbbold{1}\left(Y>\theta\right)\mid A=1,X=x\right\}-\frac{1}{t^{\prime}}\mathbb{E}\left\{\left(\theta-Y\right)\mathbbold{1}\left(\theta>Y\right)\mid A=1,X=x\right\}=0$
we have
$\frac{\widetilde{f}_{1}(\theta;t^{\prime})}{\frac{\partial}{\partial y}\theta_{1}^{+}}=t^{\prime}\mathbb{P}\left(Y>\theta\mid A=1,X=x\right)+\mathbb{P}\left(Y<\theta\mid A=1,X=x\right)=t^{\prime}\nu_{1}^{+}\{X;\theta_{1}^{+}(X;t^{\prime}),t^{\prime}\}.$
The result follows by rearranging. $\square$
Lemma 6.
Under the setup of Lemma 2,
$\frac{\partial}{\partial\theta}\mathbb{E}\{(Y-\theta)\mathbbold{1}(Y>\theta)\mid A=a,X=x\}$ $=-\mathbb{P}(Y>\theta\mid A=a,X=x)$ (42)
$\frac{\partial}{\partial\theta}\mathbb{E}\{(\theta-Y)\mathbbold{1}(\theta\geq Y)\mid A=a,X=x\}$ $=\mathbb{P}(Y\leq\theta\mid A=a,X)$ (43)
Proof:
By the continuity assumption on $p(y\mid A=a,x)$, the boundedness of $Y$, and Leibniz’ integral rule,
[…]
$=-\mathbb{P}(Y>\theta\mid A=a,X=x),$
where $\overline{B}$ is the upper bound on $Y$, and
[…]
$=\mathbb{P}(Y\leq\theta\mid A=a,X=x),$
where $\underline{B}$ is the lower bound on $Y$. $\square$
Appendix E Efficient Influence Functions
To construct doubly robust-style estimators for the bounds on the ATE we will use, when possible, the efficient influence functions (EIFs) for each part of the bound. The EIF is a concept from semiparametric efficiency theory [Bickel et al., 1993, van der Vaart, 2000, Tsiatis, 2006, van der Vaart, 2002, Van der Laan and Robins, 2003], and can be thought of as the first derivative in a von Mises expansion of an estimand [Mises, 1947]. The crucial benefit of the EIF is that estimators constructed from it are doubly robust, in the sense that they have bias that is a second-order product of errors of the relevant nuisance function estimators, so they can achieve $\sqrt{n}$-efficiency even when the nuisance functions are estimated at slower nonparametric rates [Van der Laan and Robins, 2003, Chernozhukov et al., 2018, Kennedy, 2022].
E.1 Effect differences
For the maximum LOO effects differences model (calibrated sensitivity model 1), we only require the EIF of the adjusted mean difference, provided in (14) in the body of the paper.
E.2 Odds ratio
For the maximum LOO odds ratio model (calibrated sensitivity model 2), we can use the EIF from Yadlowsky et al. [2022] for the bounds (see, e.g., (45) for the EIF of the lower bound on $\mathbb{E}(Y^{1})$). The efficient influence function for the upper bound on the ATE is
$\varphi_{U}\{Z;\eta(\Gamma M)\}$ $:=AY+(1-A)\theta_{1}^{+}\{X;\exp(\Gamma M)\}+A\frac{\omega_{\theta_{1}^{+}}(Y;\exp(\Gamma M))\pi_{0}(X)}{\nu_{1}^{+}\{X;\theta_{1}^{+},\exp(\Gamma M)\}\pi_{1}(X)}$
$-(1-A)Y+A\theta_{0}^{-}\{X;\exp(\Gamma M)\}+(1-A)\frac{\omega_{\theta_{0}^{-}}(Y;\exp(\Gamma M))\pi_{1}(X)}{\nu_{0}^{-}\{X;\theta_{1}^{+},\exp(\Gamma M)\}\pi_{0}(X)}$ (44)
Meanwhile, the efficient influence function for the lower bound on the ATE is
$\varphi_{L}\{Z;\eta(\Gamma M)\}$ $:=AY+(1-A)\theta_{1}^{-}\{X;\exp(\Gamma M)\}+A\frac{\omega_{\theta_{1}^{-}}\{Y;\exp(\Gamma M)\}\pi_{0}(X)}{\nu_{1}^{-}\{X;\theta_{1}^{-},\exp(\Gamma M)\}\pi_{1}(X)}$
$-(1-A)Y+A\theta_{0}^{+}\{X;\exp(\Gamma M)\}+(1-A)\frac{\omega_{\theta_{0}^{+}}\{Y;\exp(\Gamma M)\}\pi_{1}(X)}{\nu_{0}^{+}\{X;\theta_{1}^{+},\exp(\Gamma M)\}\pi_{0}(X)}$ (45)
where $\theta^{\pm},\nu^{\pm}$, and $\omega$ are defined in (9), (32), (33), and (25).
Because the measured confounding is a supremum, it is not pathwise differentiable and does not admit an EIF (see, e.g., Hines et al. [2022] for discussion of a lack of pathwise differentiability). Instead, we estimated its projection onto a finite dimensional model, which can be estimated at a $\sqrt{n}$-rate with a plug-in estimator, as in Definition 2.
E.3 Outcome model
The average LSO outcome model (calibrated sensitivity model 3) requires the derivation of two EIFs, for each multiplicand in the bounds in (11). The following result provides them.
Lemma 7.
The un-centered efficient influence function of $\lVert\pi_{a}(X)\rVert_{2}^{2}$ is
$\xi_{a}(Z)=\pi_{a}(X)^{2}+2\pi_{a}(X)\big{\{}\mathbbold{1}(A=a)-\pi_{a}(X)\big{\}},$ (46)
while the un-centered efficient influence function of $\lVert\mu_{a}(X_{-S})-\mathbb{E}\big{\{}\mu_{a}(X)\mid A=1-a,X_{-S}\}\rVert_{2}^{2}$ is
$\lambda_{a}(Z;S)$ $=\Big{[}\mu_{a}(X_{-S})-\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-S}\}\Big{]}^{2}$
$+2\Big{[}\mu_{a}(X_{-S})-\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-S}\}\Big{]}\Bigg{(}\frac{\mathbbold{1}(A=a)}{\pi_{a}(X_{-S})}\{Y-\mu_{a}(X_{-S})\}$
$-\left\{\frac{\mathbbold{1}(A=a)}{\pi_{a}(X)}\right\}\{Y-\mu_{a}(X)\}\left\{\frac{\pi_{1-a}(X)}{\pi_{1-a}(X_{-S})}\right\}$
$-\left\{\frac{\mathbbold{1}(A=1-a)}{\pi_{1-a}(X_{-S})}\right\}\Big{[}\mu_{a}(X)-\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-S}\}\Big{]}\Bigg{)},$ (47)
Before the proof, we provide a few comments. Each EIF follows the standard form of a plug-in plus a weighted residual. The only nuance is that the weighted residuals are multiplied by two. This occurs because each functional includes the mean of a square. Intuitively, one can think of these extra terms arising in the same way that $\frac{d}{dx}f(x^{2})=2f(x)f^{\prime}(x)$ by the chain rule. Indeed, these results are reminiscent of the EIF for the integral of the squared density, a canonical functional in the semiparametric / doubly robust estimation literature [Birgé and Massart, 1995, Laurent, 1997].
Proof:
We prove Lemma 7 in two stages. First, we show that each functional considered satisfies a von Mises expansion whose influence function is the relevant efficient influence function proposed in the result and which admits a second order remainder term. We will prove this for both functionals under consideration. Second, we will relate the von Mises expansion to smooth parametric submodels and submodel scores to prove that the proposed efficient influence function is, indeed, the efficient influence function.
Second order remainder for $\lVert\pi_{a}(X)\rVert_{2}^{2}$
We begin with $\lVert\pi_{a}(X)\rVert_{2}^{2}$. Let $\Xi_{a}(\mathcal{P})\equiv\lVert\pi_{a}(X)\rVert_{2}^{2}$ and $\xi_{a}(\mathcal{P})\equiv\xi_{a}(Z)$, where $\xi_{a}$ was defined in (46). Then, by a von Mises expansion,
$\Xi_{a}(\overline{\mathcal{P}})=\Xi_{a}(\mathcal{P})+\int_{\mathcal{Z}}\xi_{a}(\overline{\mathcal{P}})d(\overline{\mathcal{P}}-\mathcal{P})+R_{2}(\overline{\mathcal{P}},\mathcal{P})$ (48)
where $\mathcal{P}$ and $\overline{\mathcal{P}}$ are two different distributions at which $\Xi_{a}$ is evaluated. Rearranging and canceling $\Xi_{a}(\overline{\mathcal{P}})-\Xi_{a}(\overline{\mathcal{P}})$ we have
$R_{2}(\overline{\mathcal{P}},\mathcal{P})=\Xi_{a}(\overline{\mathcal{P}})-\Xi_{a}(\mathcal{P})-\int_{\mathcal{Z}}\xi_{a}(\overline{\mathcal{P}})d(\overline{\mathcal{P}}-\mathcal{P})=\int_{\mathcal{Z}}\xi_{a}(\overline{\mathcal{P}})-\xi_{a}(\mathcal{P})d(\mathcal{P})$ (49)
The right-hand side of the above equation can be re-written as an expectation over $\mathcal{P}$, but evaluated with nuisance functions from $\overline{\mathcal{P}}$ and $\mathcal{P}$; i.e.,
$R_{2}(\overline{\mathcal{P}},\mathcal{P})\equiv\mathbb{E}\left\{\overline{\xi}_{a}(Z)-\xi_{a}(Z)\right\}$
Omitting arguments, so that $\pi_{a}(X)\equiv\pi$, then, by the definition of $\xi_{a}$, we have
$R_{2}(\overline{\mathcal{P}},\mathcal{P})$ $=\mathbb{E}\Bigg{(}\overline{\pi}_{a}^{2}+2\overline{\pi}_{a}\big{\{}\mathbbold{1}(A=a)-\overline{\pi}_{a}\big{\}}-\pi_{a}^{2}-2\pi_{a}\big{\{}\mathbbold{1}(A=a)-\pi_{a}\big{\}}\Bigg{)}$
$=\mathbb{E}\left\{\overline{\pi}_{a}^{2}-\pi_{a}^{2}+2\overline{\pi}_{a}\big{(}\pi_{a}-\overline{\pi}_{a}\big{)}\right\}$
$=-\mathbb{E}\left\{\big{(}\overline{\pi}_{a}-\pi_{a}\big{)}^{2}\right\}$ (50)
where the second line follows by iterated expectations on $X$ and the third by rearranging. This shows that the remainder term is second order, which will imply that $\xi_{a}$ is the efficient influence function, as we show subsequently.
General framework for the next section
Here, we discuss an informal summary of the next section, which hopefully will help the reader as a guide as the algebra becomes more involved. First, we can informally write the efficient influence function as
$\varphi^{2}+2\rho$
so that
$R(\overline{\mathcal{P}},\mathcal{P})=\mathbb{E}\left(\overline{\varphi}^{2}+2\overline{\varphi}\overline{\rho}-\varphi^{2}-2\varphi\rho\right)$
For the section above, $\varphi$ and $\rho$ were simple enough that we could analyze the remainder term directly and show it is second order. For the next section, the algebra is more involved, but follows essentially the same format. As we will see subsequently, the algebra required to demonstrate $\lambda$ is the efficient influence functions can be split into three stages:
  1. 1.
    Show that $\mathbb{E}(\rho)=0$.
  2. 2.
    Notice then that
    $\mathbb{E}\left(\overline{\varphi}^{2}+2\overline{\rho}-\varphi^{2}-2\rho\right)=\mathbb{E}\left\{-(\overline{\varphi}-\varphi)^{2}+2(\overline{\rho}+\overline{\varphi}^{2}-\overline{\varphi}\varphi)\right\}$
    using 1. and then completing the square.
  3. 3.
    Show that $-\mathbb{E}\{(\overline{\varphi}-\varphi)^{2}\}$ is second order.
  4. 4.
    Show that $\mathbb{E}\{\overline{\rho}+\overline{\varphi}^{2}-\overline{\varphi}\varphi\}$ is second order.
Second order remainder for $\lVert\mu_{a}(X_{-S})-\mathbb{E}\big{\{}\mu_{a}(X)\mid A=1-a,X_{-S}\}\rVert_{2}^{2}$
Here, we consider $\lVert\mu_{a}(X_{-S})-\mathbb{E}\big{\{}\mu_{a}(X)\mid A=1-a,X_{-S}\}\rVert_{2}^{2}$. By the same logic as above,
$R_{2}(\overline{\mathcal{P}},\mathcal{P})=\mathbb{E}\{\overline{\lambda}_{a}(Z;S)-\lambda_{a}(Z;S)\}.$ (51)
Let
$\varphi_{a}(X_{-S})$ $=\mu_{a}(X_{-S})-\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-S}\}\text{ and }$
$\rho_{a}(Z;S)$ $=\Big{[}\mu_{a}(X_{-S})-\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-S}\}\Big{]}\Bigg{(}\frac{\mathbbold{1}(A=a)}{\pi_{a}(X_{-S})}\{Y-\mu_{a}(X_{-S})\}$
$-\left\{\frac{\mathbbold{1}(A=a)}{\pi_{a}(X)}\right\}\{Y-\mu_{a}(X)\}\left\{\frac{\pi_{1-a}(X)}{\pi_{1-a}(X_{-S})}\right\}$
$-\left\{\frac{\mathbbold{1}(A=1-a)}{\pi_{1-a}(X_{-S})}\right\}\Big{[}\mu_{a}(X)-\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-S}\}\Big{]}\Bigg{)}$
so that
$\lambda_{a}(Z;S)=\varphi_{a}(X_{-S})^{2}+2\rho_{a}(Z;S).$
Notice that
$\mathbb{E}\left\{\rho_{a}(Z;S)\right\}=0$ (52)
by iterated expectations on $A=a,X$ and then iterated expectations on $A=a,X_{-S}$ and $A=1-a,X_{-S}$. Second, as in the informal proof sketch, notice that
$R_{2}(\overline{\mathcal{P}},\mathcal{P})=\mathbb{E}\left[-\{\overline{\varphi}_{a}(X_{-S})-\varphi_{a}(X_{-S})\}^{2}+2\Big{\{}\overline{\varphi}_{a}(X_{-S})^{2}+\overline{\rho}_{a}(Z;S)-\overline{\varphi}_{a}(X_{-S})\varphi_{a}(X_{-S})\}\right]$ (53)
which follows by (51), the definition of $\lambda,\varphi,$ and $\rho$, (52), and completing the square. Third, notice that $\mathbb{E}\left[\{\overline{\varphi}_{a}(X_{-S})-\varphi_{a}(X_{-S})\}^{2}\right]$ is second order in terms of $\mu_{a}(X_{-S})$ and $\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-S}\}$. Fourth, notice that
$\mathbb{E}\big{\{}\overline{\rho}_{a}(Z;S)\}$ $=\mathbb{E}\Bigg{\{}\Big{[}\overline{\mu}_{a}(X_{-S})-\overline{\mathbb{E}}\{\overline{\mu}_{a}(X)\mid A=1-a,X_{-S}\}\Big{]}\Bigg{(}\frac{\mathbbold{1}(A=a)}{\overline{\pi}_{a}(X_{-S})}\{Y-\overline{\mu}_{a}(X_{-S})\}$
$-\left\{\frac{\mathbbold{1}(A=a)}{\overline{\pi}_{a}(X)}\right\}\{Y-\overline{\mu}_{a}(X)\}\left\{\frac{\overline{\pi}_{1-a}(X)}{\overline{\pi}_{1-a}(X_{-S})}\right\}$
$-\left\{\frac{\mathbbold{1}(A=1-a)}{\overline{\pi}_{1-a}(X_{-S})}\right\}\Big{[}\overline{\mu}_{a}(X)-\overline{\mathbb{E}}\{\overline{\mu}_{a}(X)\mid A=1-a,X_{-S}\}\Big{]}\Bigg{)}\Bigg{\}}$
$=\mathbb{E}\Bigg{\{}\Big{[}\overline{\mu}_{a}(X_{-S})-\overline{\mathbb{E}}\{\overline{\mu}_{a}(X)\mid A=1-a,X_{-S}\}\Big{]}\Bigg{(}\frac{\pi_{a}(X_{-S})}{\overline{\pi}_{a}(X_{-S})}\{\mu_{a}(X_{-S})-\overline{\mu}_{a}(X_{-S})\}$
$-\left\{\frac{\pi_{a}(X)}{\overline{\pi}_{a}(X)}\right\}\{\mu_{a}(X)-\overline{\mu}_{a}(X)\}\left\{\frac{\overline{\pi}_{1-a}(X)}{\overline{\pi}_{1-a}(X_{-S})}\right\}$
$-\left\{\frac{\pi_{1-a}(X_{-S})}{\overline{\pi}_{1-a}(X_{-S})}\right\}\Big{[}(\mathbb{E}-\overline{\mathbb{E}})\{\overline{\mu}_{a}(X)\mid A=1-a,X_{-S}\}\Big{]}\Bigg{)}\Bigg{\}}$
$=\mathbb{E}\Bigg{\{}\overline{\varphi}_{a}(X_{-S})\Bigg{(}\left\{\frac{\pi_{a}(X_{-S})-\overline{\pi}_{a}(X_{-S})}{\overline{\pi}_{a}(X_{-S})}\right\}\{\mu_{a}(X_{-S})-\overline{\mu}_{a}(X_{-S})\}+\{\mu_{a}(X_{-S})-\overline{\mu}_{a}(X_{-S})\}$
$-\left\{\frac{\pi_{a}(X)-\overline{\pi}_{a}(X}{\overline{\pi}_{a}(X)}\right\}\{\mu_{a}(X)-\overline{\mu}_{a}(X)\}\left\{\frac{\overline{\pi}_{1-a}(X)}{\overline{\pi}_{1-a}(X_{-S})}\right\}$
$-\{\mu_{a}(X_{[p})-\overline{\mu}_{a}(X)\}\left\{\frac{\overline{\pi}_{1-a}(X)}{\overline{\pi}_{1-a}(X_{-S})}\right\}$
$-\left\{\frac{\pi_{1-a}(X_{-S})-\overline{\pi}_{1-a}(X_{-S})}{\overline{\pi}_{1-a}(X_{-S})}\right\}\Big{[}(\mathbb{E}-\overline{\mathbb{E}})\{\overline{\mu}_{a}(X)\mid A=1-a,X_{-S}\}\Big{]}$
$-(\mathbb{E}-\overline{\mathbb{E}})\{\overline{\mu}_{a}(X)\mid A=1-a,X_{-S}\}\Bigg{)}\Bigg{\}}$
$=\mathbb{E}\Bigg{\{}\overline{\varphi}_{a}(X_{-S})\Bigg{(}\left\{\frac{\pi_{a}(X_{-S})-\overline{\pi}_{a}(X_{-S})}{\overline{\pi}_{a}(X_{-S})}\right\}\{\mu_{a}(X_{-S})-\overline{\mu}_{a}(X_{-S})\}$
$-\left\{\frac{\pi_{a}(X)-\overline{\pi}_{a}(X)}{\overline{\pi}_{a}(X)}\right\}\{\mu_{a}(X)-\overline{\mu}_{a}(X)\}\left\{\frac{\overline{\pi}_{1-a}(X)}{\overline{\pi}_{1-a}(X_{-S})}\right\}$
$-\left\{\frac{\pi_{1-a}(X_{-S})-\overline{\pi}_{1-a}(X_{-S})}{\overline{\pi}_{1-a}(X_{-S})}\right\}\Big{[}(\mathbb{E}-\overline{\mathbb{E}})\{\overline{\mu}_{a}(X)\mid A=1-a,X_{-S}\}\Big{]}$
$-\{\mu_{a}(X_{[p})-\overline{\mu}_{a}(X)\}\left\{\frac{\overline{\pi}_{1-a}(X)}{\overline{\pi}_{1-a}(X_{-S})}-\frac{\pi_{1-a}(X)}{\pi_{1-a}(X_{-S})}\right\}$
$-\{\mu_{a}(X_{[p})-\overline{\mu}_{a}(X)\}\left\{\frac{\pi_{1-a}(X)}{\pi_{1-a}(X_{-S})}\right\}$
$-(\mathbb{E}-\overline{\mathbb{E}})\{\overline{\mu}_{a}(X)\mid A=1-a,X_{-S}\}+\{\mu_{a}(X_{-S})-\overline{\mu}_{a}(X_{-S})\}\Bigg{)}\Bigg{\}}$
where the first line follows definition, the second by iterated expectations on $X$ and $A=a$ and then on $X_{-S}$, the third by adding zero, and the fourth again by adding zero and rearranging. Notice that $\overline{\varphi}$ was factored out at the start and that the first four summands are second order. Focusing on the last two summands, notice that
$\mathbb{E}\Bigg{(}\overline{\varphi}_{a}(X_{-S})$ [$\Bigg{[}-\{\mu_{a}(X_{[p})-\overline{\mu}_{a}(X)\}\left\{\frac{\pi_{1-a}(X)}{\pi_{1-a}(X_{-S})}\right\}$
$-(\mathbb{E}-\overline{\mathbb{E}})\{\overline{\mu}_{a}(X)\mid A=1-a,X_{-S}\}+\{\mu_{a}(X_{-S})-\overline{\mu}_{a}(X_{-S})\}\Bigg{]}\Bigg{)}$
$=\mathbb{E}\Bigg{(}\overline{\varphi}_{a}(X_{-S})\Bigg{[}$$-\{\mu_{a}(X_{[p})-\overline{\mu}_{a}(X)\}\left\{\frac{\pi_{1-a}(X)}{\pi_{1-a}(X_{-S})}\right\}$
$-\overline{\varphi}(X_{-S})+\mu_{a}(X_{-S})-\mathbb{E}\{\overline{\mu}_{a}(X)\mid 1-a,X_{-S}\}]$
$=\mathbb{E}\Bigg{[}\overline{\varphi}_{a}(X_{-S})\Bigg{\{}\varphi_{a}(X_{-S})-\overline{\varphi}_{a}(X_{-S})\Bigg{\}}\Bigg{]}$
where the second line follows by the definition of $\overline{\varphi}_{a}(X_{-S})$ and the final line follows because
[…]
$=\int_{\mathcal{X}_{-S}}\{\mu_{a}(X_{[p})-\overline{\mu}_{a}(x)\}\mathbb{P}(x\mid A=1-a,X_{-S})$
$=\mathbb{E}\{\mu_{a}(X)-\overline{\mu}_{a}(X)\mid A=1-a,X_{-S}\}$
where the first line follows by definition, the second because, under Assumption 2, $\frac{\pi_{1-a}(X)}{\pi_{1-a}(X_{-S})}=\frac{\mathbb{P}(X\mid A-1,X_{-S})}{\mathbb{P}(X\mid X_{-S})}$, the third by canceling, and the final line by definition.
Therefore, we have
$\mathbb{E}\{\rho_{a}(Z;S)\}=\mathbb{E}\Big{[}f(Z;S)-\overline{\varphi}_{a}(X_{-S})^{2}+\overline{\varphi}_{a}(X_{-S})\varphi_{a}(X_{-S})\Big{]}$
where $\mathbb{E}\{f(Z;S)\}$ is a second order product of errors. Plugging this back into (53), notice that
$R_{2}(\overline{\mathcal{P}},\mathcal{P})=\mathbb{E}\left[-\{\overline{\varphi}_{a}(X_{-S})-\varphi_{a}(X_{-S})\}^{2}+2\Big{\{}f(Z;S)\Big{\}}\right]$ (54)
where $\mathbb{E}\{f(Z;S)\}$ is the second order product of errors; specifically, we have
$\mathbb{E}\{f(Z;S)\}=\mathbb{E}\Bigg{\{}\overline{\varphi}_{a}(X_{-S})\Bigg{(}$$\left\{\frac{\pi_{a}(X_{-S})-\overline{\pi}_{a}(X_{-S})}{\overline{\pi}_{a}(X_{-S})}\right\}\{\mu_{a}(X_{-S})-\overline{\mu}_{a}(X_{-S})\}$
$-\left\{\frac{\pi_{a}(X)-\overline{\pi}_{a}(X)}{\overline{\pi}_{a}(X)}\right\}\{\mu_{a}(X)-\overline{\mu}_{a}(X)\}\left\{\frac{\overline{\pi}_{1-a}(X)}{\overline{\pi}_{1-a}(X_{-S})}\right\}$
$-\left\{\frac{\pi_{1-a}(X_{-S})-\overline{\pi}_{1-a}(X_{-S})}{\overline{\pi}_{1-a}(X_{-S})}\right\}\Big{[}(\mathbb{E}-\overline{\mathbb{E}})\{\overline{\mu}_{a}(X)\mid A=1-a,X_{-S}\}\Big{]}$
$-\{\mu_{a}(X_{[p})-\overline{\mu}_{a}(X)\}\left\{\frac{\overline{\pi}_{1-a}(X)}{\overline{\pi}_{1-a}(X_{-S})}-\frac{\pi_{1-a}(X)}{\pi_{1-a}(X_{-S})}\right\}\Bigg{)}\Bigg{\}}.$ (55)
Therefore, we conclude that the remainder term is second-order.
Proof that second order remainder implies efficient influence function
Now, we relate each of the efficient influence functions back to scores of smooth parametric submodels. We will make a general argument that applies to both efficient influence functions. We ignore subscript notation, and let $\psi:\mathcal{P}\to\mathbb{R}$ denote a generic functional. We include this last part of the proof for completeness, but it is an imitation of Lemma 2 in Kennedy et al. [2023].
Recall from semiparametric efficiency theory that the nonparametric efficiency bound for a functional is given by the supremum of Cramer-Rao lower bounds for that functional across smooth parametric submodels [Bickel et al., 1993, van der Vaart, 2002]. The efficient influence function is the unique mean-zero function that is a valid submodel score satisfying pathwise differentiability; i.e., if $\varphi$ is the efficient influence function of $\psi$, it satisfies
$\frac{d}{d\epsilon}\psi(\mathcal{P}_{\epsilon})\bigg{|}_{\epsilon=0}=\int_{\mathcal{Z}}\varphi(\mathcal{P})\left(\frac{d}{d\epsilon}\log d\mathcal{P}_{\epsilon}\right)\bigg{|}_{\epsilon=0}d\mathcal{P}$ (56)
for $\mathcal{P}_{\epsilon}$ any smooth parametric submodel.
To see that each of the proposed efficient influence functions are, indeed, efficient influence functions, observe that the relevant von Mises expansions imply
$\frac{d}{d\epsilon}\psi(\mathcal{P}_{\epsilon})$ $=\frac{d}{d\epsilon}\left(\psi(\mathcal{P})-\int_{\mathcal{Z}}\varphi(\mathcal{P})d(\mathcal{P}-\mathcal{P}_{\epsilon})-R_{2}(\mathcal{P},\mathcal{P}_{\epsilon})\right)$
$=\frac{d}{d\epsilon}\int_{\mathcal{Z}}\varphi(\mathcal{P})d(\mathcal{P}_{\epsilon}-\mathcal{P})-\frac{d}{d\epsilon}R_{2}(\mathcal{P},\mathcal{P}_{\epsilon})$
[…]
since
$\frac{d}{d\epsilon}R_{2}(\mathcal{P},\mathcal{P}_{\epsilon})\Big{|}_{\epsilon=0}=0,$
which shows that $\varphi$ satisfies the property in eq. (56). The last equation involving $R_{2}$ follows because $R_{2}$ consists of only second-order products of errors between $\mathcal{P}_{\epsilon}$ and $\mathcal{P}$ by the arguments above. Therefore, the derivative is composed of a sum of terms, each of which is a product of a derivative term that may not equal zero and an error term involving the differences of components of $\mathcal{P}$ and $\mathcal{P}_{\epsilon}$, which will be zero when $\epsilon=0$ since $\mathcal{P}=\mathcal{P}_{\epsilon}$.
Since the model is nonparametric, the tangent space is the entire Hilbert space of mean-zero finite-variance functions, and so there is only one influence function satisfying eq. (56) and it is the efficient one [Tsiatis, 2006]. Therefore, both efficient influence functions proposed in Lemma 7 are indeed, the efficient influence functions. $\square$
Appendix F Proofs for Section 5
Here, we prove Theorems 1-3. Each proof is split into two steps:
  1. 1.
    Establishing that the estimator for measured confounding is regular and asymptotically linear (possibly under doubly robust conditions), and
  2. 2.
    Establishing that the estimator for the upper bound on the ATE is regular and asymptotically linear.
Let $\mathbb{P}$ denote expectation conditional on the training sample throughout what follows. Moreover, note that the positivity assumption (Assumption 2) implies $\mu_{a}(X_{-S})$ are well-defined for all $S\in 2^{[d]}$.
Proof of Theorem 1
Proof:
Here, measured confounding is $|\psi-\psi_{-j^{\prime}}|$. Therefore, first, we establish convergence guarantees for each adjusted mean difference estimator. For all $j\in[d]\cup\emptyset$, we have
$\widehat{\psi}_{-j}-\psi_{-j}=(\mathbb{P}_{n}-\mathbb{E})\{\phi(Z_{-j})\}+(\mathbb{P}_{n}-\mathbb{P})\{\widehat{\phi}(Z_{-j})-\phi(Z_{-j})\}+\mathbb{P}\{\widehat{\phi}(Z_{-j})-\phi(Z_{-j})\}$ (57)
by adding zero, and where $\phi$ is defined in (14). Then, by condition 2 of the theorem, sample splitting, and Lemma 2 in Kennedy et al. [2020], the empirical process term satisfies $(\mathbb{P}_{n}-\mathbb{P})\{\widehat{\phi}(Z_{-j})-\phi(Z_{-j})\}=o_{\mathbb{P}}(n^{-1/2})$. Moreover, by condition 2 and Jensen’s inequality, the bias satisfies
$\mathbb{P}\{\widehat{\phi}(Z_{-j})-\phi(Z_{-j})\}\leq\sqrt{\mathbb{P}\Big{[}\{\widehat{\phi}(Z_{-j})-\phi(Z_{-j})\}^{2}\Big{]}}\equiv\lVert\widehat{\phi}(Z_{-j})-\phi(Z_{-j})\rVert_{2}=o_{\mathbb{P}}(1).$
Hence, by applying the standard CLT to $(\mathbb{P}_{n}-\mathbb{E})\{\phi(Z_{-j})\}$,
$\widehat{\psi}_{-j}-\psi_{-j}=o_{\mathbb{P}}(1)\ \forall\ j\in[d]\cup\emptyset.$
For $j\in\{j^{\prime},\emptyset\}$, the convergence guarantee can be strengthened. We have
$\widehat{\psi}_{-j}-\psi_{-j}=(\mathbb{P}_{n}-\mathbb{E})\{\phi(Z_{-S})\}+o_{\mathbb{P}}(n^{-1/2})$
because the bias term from (57) satisfies $\mathbb{P}\{\widehat{\phi}(Z_{-j})-\phi(Z_{-j})\}=o_{\mathbb{P}}(n^{-1/2})$ by iterated expectations, Cauchy-Schwarz, and conditions 1 and 3 of the theorem.
Next, we can analyze the convergence of the estimator for measured confounding, $\max_{j\in[d]}|\widehat{\psi}-\widehat{\psi}_{-j}|$. Because $f(x,y)=|x-y|$ is continuous for all $x,y\in\mathbb{R}$,
$|\widehat{\psi}-\widehat{\psi}_{-j}|-|\psi-\psi_{-j}|=o_{\mathbb{P}}(1)\ \forall\ j\in[d]$ (58)
by the continuous mapping theorem. Moreover, by Assumption 4, which guarantees $|\psi-\psi_{-j^{\prime}}|$ is unique, and the delta method,
$|\widehat{\psi}-\widehat{\psi}_{-j^{\prime}}|-|\psi-\psi_{-j^{\prime}}|=(\mathbb{P}_{n}-\mathbb{E})\big{[}\text{sign}\left(\psi-\psi_{-j^{\prime}}\right)\left\{\phi(Z)-\phi(Z_{-j^{\prime}})\right\}\big{]}+o_{\mathbb{P}}(n^{-1/2}).$ (59)
Therefore, by Theorem 4 and (58) and (59),
$\max_{j\in[d]}|\widehat{\psi}-\widehat{\psi}_{-j}|-|\psi-\psi_{j^{\prime}}|=(\mathbb{P}_{n}-\mathbb{E})\big{[}\text{sign}\left(\psi-\psi_{-j^{\prime}}\right)\left\{\phi(Z)-\phi(Z_{-j^{\prime}})\right\}\big{]}+o_{\mathbb{P}}(n^{-1/2}).$
Finally, because $\widehat{\psi}-\psi=(\mathbb{P}_{n}-\mathbb{E})\{\phi(Z)\}+o_{\mathbb{P}}(n^{-1/2})$, the result then follows by the delta method and because $0<\Gamma|\psi-\psi_{j^{\prime}}|<\infty$ by Assumption 3. $\square$
Proof of Theorem 2
Proof:
We begin with the estimator for measured confounding. By conditions 1 and 2 of the result, $\widehat{\pi}_{1}^{\perp}$ is a regular and asymptotically linear estimator for $\pi_{1}^{\perp}$ by Theorem 5.39 in van der Vaart [2000]; moreover, the estimator for the $j^{th}$ coefficient of $\beta$ satisfies
$\widehat{\beta}_{j}-\beta_{j}=e_{j}^{T}I_{\beta}^{-1}(\mathbb{P}_{n}-\mathbb{E})\big{\{}s(Z;\beta)\big{\}}+o_{\mathbb{P}}(n^{-1/2})$
by the delta method, where $I$ and $s$ are defined in the result. Then, as with the proof of Theorem 1, by the continuous mapping theorem, delta method, and Theorem 4,
$\widehat{M}-M=(\mathbb{P}_{n}-\mathbb{P})\{\phi(Z)\}+o_{\mathbb{P}}(n^{-1/2}),$
where $\phi_{M}(Z)$ is defined in the result, and $\widehat{M}$ and $M$ are given in Definition 2.
Next, we analyze the estimator for the bound. By iterated expectations in the first line and adding zero in the second line,
$\mathbb{P}_{n}\big{[}\varphi_{\mathcal{U}}\{Z;\widehat{\eta}(\Gamma\widehat{M})\}\big{]}-\mathcal{U}(\Gamma)$ $=\mathbb{P}_{n}\big{[}\varphi_{\mathcal{U}}\{Z;\widehat{\eta}(\Gamma\widehat{M})\}\big{]}-\mathbb{E}\big{[}\varphi_{\mathcal{U}}\{Z;\eta(\Gamma M)\}\big{]}$
$=(\mathbb{P}_{n}-\mathbb{E})\big{[}\varphi_{\mathcal{U}}\{Z;\eta(\Gamma M)\}\big{]}$
$+(\mathbb{P}_{n}-\mathbb{P})\big{[}\varphi_{U}\{Z;\widehat{\eta}(\Gamma\widehat{M})\}-\varphi_{U}\{Z;\eta(\Gamma M)\}\big{]}$
$+\mathbb{P}\big{[}\varphi_{U}\{Z;\widehat{\eta}(\Gamma\widehat{M})\}-\varphi_{U}\{Z;\eta(\Gamma\widehat{M})\}\big{]}$
$+\mathbb{P}\big{[}\varphi_{U}\{Z;\eta(\Gamma\widehat{M})\}-\varphi_{U}\{Z;\eta(\Gamma M)\}\big{]},$
where $\varphi_{\mathcal{U}}$ is defined in (44). The empirical process term $(\mathbb{P}_{n}-\mathbb{P})\big{[}\varphi_{U}\{Z;\widehat{\eta}(\Gamma\widehat{M})\}-\varphi_{U}\{Z;\eta(\Gamma M)\}\big{]}=o_{\mathbb{P}}(n^{-1/2})$ by sample splitting, Lemma 2 in Kennedy et al. [2020], and condition 3 of the result. The first bias term satisfies $\mathbb{P}\big{[}\varphi_{U}\{Z;\widehat{\eta}(\Gamma\widehat{M})\}-\varphi_{U}\{Z;\eta(\Gamma\widehat{M})\}\big{]}=o_{\mathbb{P}}(n^{-1/2})$ by condition 4 of the result. The second bias term can be decomposed using Lemma 2 and Taylor’s theorem:
$\mathbb{P}\big{[}\varphi_{U}\{Z;\eta(\Gamma\widehat{M})\}-\varphi_{U}\{Z;\eta(\Gamma M)\}\big{]}=\frac{\partial}{\partial t}\mathbb{E}\big{[}\varphi_{U}\{Z;\eta(\Gamma t)\}\big{]}\Big{|}_{t=M}\left(\Gamma\widehat{M}-\Gamma M\right)+o\left(\Gamma\left|\widehat{M}-M\right|\right).$
Notice that, by iterated expectations, $\frac{\partial}{\partial t}\mathbb{E}\big{[}\varphi_{U}\{Z;\eta(\Gamma t)\}\big{]}\Big{|}_{t=M}\equiv\frac{\partial}{\partial M}\mathcal{U}(\Gamma)$, where $\frac{\partial}{\partial M}\mathcal{U}(\Gamma)$ is defined in Lemma 1. Meanwhile, $o\left(\Gamma\left|\widehat{M}-M\right|\right)=o_{\mathbb{P}}(n^{-1/2})$ because $\widehat{M}-M=O_{\mathbb{P}}(n^{-1/2})$ by the argument above for estimating measured confounding. The result then follows by the delta method. $\square$
Proof of Theorem 3
Proof:
First, we consider estimating the components of measured confounding. To begin, notice that
$\mathbb{P}_{n}\{\widehat{\xi}_{a}(X)\}-\lVert\pi_{a}(X)\rVert_{2}^{2}=(\mathbb{P}_{n}-\mathbb{E})\{\xi_{a}(X)\}+(\mathbb{P}_{n}-\mathbb{P})\{\widehat{\xi}_{a}(X)-\xi_{a}(X)\}+\mathbb{P}\{\widehat{\xi}_{a}(X)-\xi_{a}(X)\},$
where $\xi$ is defined in (46). Then, by condition 3 of the result, sample splitting, and Lemma 2 of Kennedy et al. [2020], the empirical process term satisfies $(\mathbb{P}_{n}-\mathbb{P})\{\widehat{\xi}_{a}(X)-\xi_{a}(X)\}=o_{\mathbb{P}}(n^{-1/2})$. Next, notice that the bias term is a second-order product of errors by the same proof as for deriving the EIF in Lemma 7. Indeed, the bias is second-order by (50), but replacing overlines by hats. Therefore, by condition 6 and Cauchy-Schwarz, $\mathbb{P}\{\widehat{\xi}_{a}(X)-\xi_{a}(X)\}=o_{\mathbb{P}}(n^{-1/2}).$ Therefore,
$\mathbb{P}_{n}\{\widehat{\xi}_{a}(X)\}-\lVert\pi_{a}(X)\rVert_{2}^{2}=(\mathbb{P}_{n}-\mathbb{E})\{\xi_{a}(X)\}+o_{\mathbb{P}}(n^{-1/2}).$
So, by Assumption 2 and the delta method,
$\sqrt{\mathbb{P}_{n}\{\widehat{\xi}_{a}(X)\}}-\lVert\pi_{a}(X)\rVert_{2}=(\mathbb{P}_{n}-\mathbb{E})\left\{\frac{\xi_{a}(X)}{2\lVert\pi_{a}(X)\rVert_{2}}\right\}+o_{\mathbb{P}}(n^{-1/2}).$
Next,
$\mathbb{P}_{n}\{\widehat{\lambda}_{a}(Z;S)\}-\lVert\mu_{a}(X_{-S})-\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-S}\}\rVert_{2}^{2}$
$\hskip 36.135pt=(\mathbb{P}_{n}-\mathbb{E})\{\lambda_{a}(Z;S)\}+(\mathbb{P}_{n}-\mathbb{P})\{\lambda_{a}(Z;S)-\lambda_{a}(Z;S)\}+\mathbb{P}\{\widehat{\lambda}_{a}(Z;S)-\lambda_{a}(Z;S)\},$
where $\lambda$ is defined in (47). The empirical process term is $o_{\mathbb{P}}(n^{-1/2})$ by sample splitting, Lemma 2 of Kennedy et al. [2020], and condition 4. Then, (54) and (55) define the bias of $\mathbb{P}\{\widehat{\lambda}_{a}(Z;S)-\lambda_{a}(Z;S)\}$, replacing overlines by hats. Therefore, the bias term can be controlled by conditions 1, 5, and 7 and Cauchy-Schwarz. In conclusion,
$\mathbb{P}_{n}\{\widehat{\lambda}_{a}(Z;S)\}-\lVert\mu_{a}(X_{-S})-\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-S}\}\rVert_{2}^{2}=(\mathbb{P}_{n}-\mathbb{E})\{\lambda_{a}(Z;S)\}+o_{\mathbb{P}}(n^{-1/2})$
for $a\in\{0,1\}$ and all $S\in\mathcal{S}$. Then, because measured confounding is non-zero by Assumption 3, the delta method yields
$\frac{1}{|\mathcal{S}|}\sum_{S\in\mathcal{S}}\mathbb{P}_{n}\{\widehat{\lambda}_{a}(Z;S)\}-\frac{1}{|\mathcal{S}|}\sum_{S\in\mathcal{S}}\lVert\mu_{a}(X_{-S})-\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-s}\}\rVert_{2}^{2}=\frac{1}{|\mathcal{S}|}\sum_{S\in\mathcal{S}}(\mathbb{P}_{n}-\mathbb{P})\{\lambda_{a}(Z;S)\}+o_{\mathbb{P}}(n^{-1/2})$
and
$\sqrt{\frac{1}{|\mathcal{S}|}\sum_{S\in\mathcal{S}}\mathbb{P}_{n}\{\widehat{\lambda}_{a}(Z;S)\}}-\sqrt{\frac{1}{|\mathcal{S}|}\sum_{S\in\mathcal{S}}\lVert\mu_{a}(X_{-S})-\mathbb{E}\{\mu_{a}(X)\mid A=1-a,X_{-s}\}\rVert_{2}^{2}}=\frac{1}{|\mathcal{S}|}\sum_{S\in\mathcal{S}}(\mathbb{P}_{n}-\mathbb{P})\left\{\frac{\lambda_{a}(Z;S)}{2M_{a}}\right\}+o_{\mathbb{P}}(n^{-1/2})$
where $M_{a}$ is defined in the result.
Finally, for the other piece of the bound, $\mathbb{P}_{n}\{\widehat{\phi}(Z)\}-\psi=(\mathbb{P}_{n}-\mathbb{E})\{\phi(Z)\}+o_{\mathbb{P}}(n^{-1/2})$, by conditions 1, 2, and 5, and the same analysis as in the proof of Theorem 1. The result then follows by the delta method. $\square$
Appendix G Technical results for the estimator of maximum measured confounding
Here, we state and prove a helper result that was used in the proofs of Theorems 1 and 2, which, in context of this paper, implies that the estimator of maximum measured confounding is regular and asymptotically linear providing the estimator of the measured confounding corresponding to the maximum is regular and asymptotically linear and the estimators for all the non-maximum sub-quantifications of measured confounding are consistent.
Theorem 4.
Suppose, for $J<\infty$ bounded real-valued functionals $\psi_{1},\dots,\psi_{J}$,
  1. 1.
    without loss of generality, $\psi_{1}>\psi_{j}$ for all $j>1$,
  2. 2.
    $\widehat{\psi}_{1}$ satisfies $\widehat{\psi}_{1}-\psi_{1}=(\mathbb{P}_{n}-\mathbb{P})\varphi_{1}+o_{\mathbb{P}}(n^{-1/2})$, and
  3. 3.
    $\widehat{\psi}_{j}$ satisfies $\widehat{\psi}_{j}-\psi_{j}=o_{\mathbb{P}}(1)$ for all $j>1$.
Let
$\widehat{j}=\operatorname*{arg\,max}_{j\in\{1,\dots,J\}}\widehat{\psi}_{j}.$ (60)
Then,
$\widehat{\psi}_{\widehat{j}}-\psi_{1}=(\mathbb{P}_{n}-\mathbb{P})\varphi_{1}+o_{\mathbb{P}}(n^{-1/2})$ (61)
Proof:
We can decompose the error like so:
$\widehat{\psi}_{\widehat{j}}-\psi_{1}=\widehat{\psi}_{\widehat{j}}-\widehat{\psi}_{1}+\widehat{\psi}_{1}-\psi_{1}=\underbrace{\mathbbold{1}(\widehat{j}\neq 1)\left(\widehat{\psi}_{\widehat{j}}-\widehat{\psi}_{1}\right)}_{R_{n}}+\underbrace{\widehat{\psi}_{1}-\psi_{1}}_{Z}.$
The second term $Z$ satisfies the result. Meanwhile, $\mathbbold{1}(\widehat{j}\neq 1)=o_{\mathbb{P}}(n^{-1/2})$ by Lemma 8. Finally, notice that
$\left|\widehat{\psi}_{\widehat{j}}-\widehat{\psi}_{1}\right|\leq\left|\widehat{\psi}_{\widehat{j}}-\psi_{\widehat{j}}\right|+\left|\psi_{\widehat{j}}-\psi_{1}\right|+\left|\psi_{1}-\widehat{\psi}_{1}\right|=O_{\mathbb{P}}(1),$
where the last equality follows by conditions 2 and 3 and because $|\psi_{\widehat{j}}-\psi_{1}|$ is bounded by assumption. Therefore, because $|X|=O_{\mathbb{P}}(1)\implies X=O_{\mathbb{P}}(1)$ for any random variable $X$,
$R_{n}=o_{\mathbb{P}}(n^{-1/2})O_{\mathbb{P}}(1)=o_{\mathbb{P}}(n^{-1/2}).$
The result follows. $\square$
Lemma 8.
Under the setup of Theorem 4, but weakening condition 2 such that $\widehat{\psi}_{1}-\psi_{1}=o_{\mathbb{P}}(1)$ (i.e., such that $\widehat{\psi}_{1}$ is merely consistent),
$\mathbbold{1}(\widehat{j}\neq 1)=o_{\mathbb{P}}(n^{-\alpha})$
for all $\alpha>0$.
Proof:
Let $\epsilon>0$. For all $n>\epsilon^{1/\alpha}$,
$\mathbb{P}\left\{|n^{\alpha}\mathbbold{1}(\widehat{j}\neq 1)|>\epsilon\right\}$ $=\mathbb{P}(\widehat{j}\neq 1)$
$\leq\sum_{j\in\{2,\dots,J\}}\mathbb{P}\left(\widehat{\psi}_{1}<\widehat{\psi}_{j}\right),$
where the inequality follows by the union bound. Notice that the final term on the right-hand side converges to zero as $n\to\infty$ because, for all $j>1$,
$\mathbb{P}(\widehat{\psi}_{1}<\widehat{\psi}_{j})$ $=\mathbb{P}\left(\widehat{\psi}_{j}-\psi_{j}+\psi_{1}-\widehat{\psi}_{1}>\psi_{1}-\psi_{j}\right)$
$\leq\mathbb{P}\left(|\widehat{\psi}_{j}-\psi_{j}|+|\widehat{\psi}_{1}-\psi_{1}|>\psi_{1}-\psi_{j}\right)$
$\leq\mathbb{P}\left(|\widehat{\psi}_{j}-\psi_{j}|>\frac{\psi_{1}-\psi_{j}}{2}\right)+\mathbb{P}\left(|\widehat{\psi}_{1}-\psi_{1}|>\frac{\psi_{1}-\psi_{j}}{2}\right)$
$\to 0\text{ as }n\to\infty,$
where the first line follows by adding zero and rearranging, the third by the union bound and because, for $a,b,c>0$, $a+b>c\implies a>c/2\text{ or }b>c/2$, and the fourth because $\widehat{\psi}_{j}\stackrel{{\scriptstyle p}}{{\rightarrow}}\psi_{j}$ for all $j\in[J]$ by conditions 2 and 3 of Theorem 4 and because $\psi_{1}>\psi_{j}$ for all $j>1$ by condition 1 of Theorem 4. Hence, $\mathbbold{1}(\widehat{j}\neq 1)=o_{\mathbb{P}}(n^{-\alpha})$ by the definition of $o_{\mathbb{P}}$ notation. $\square$